### v2

# arXiv:2509.24762v2[cs.LG] 27 Feb 2026

Abstract
Modeling multi-type event sequences with marked temporal point processes (MTPPs) provides a principled framework for uncovering governing dynamical rules and predicting future events. Current neural approaches to MTPP inference typically require training separate, specialized models for each target system. We pursue a fundamentally different strategy: leveraging amortized inference and in-context learning, we pretrain a deep neural network to infer, in-context, the conditional intensity functions of event histories from a context consisting of sets of event sequences. Pretraining is performed on a large synthetic dataset of MTPPs sampled from a broad distribution over point processes. Once pretrained, our Foundation Inference Model for Point Processes (FIM-PP) can estimate MTPPs from real-world data without additional training, or be rapidly finetuned to specific target systems. Experiments show that FIM-PP matches the performance of specialized models on multi-event prediction across common benchmark datasets.
Our pretrained model, repository, and tutorials are available online [^1] https://fim4science.github.io/OpenFIM/intro.html
1 Introduction
The mathematical modeling of asynchronous, irregular event sequences has long occupied a distinctive place in the machine learning community. Temporal point processes provide the canonical framework for modeling neural dynamics (Truccolo et al., 2005; Linderman and Adams, 2014), and they serve as the de facto tool for describing a wide range of internet phenomena, including retweeting, posting, and information cascades (Zhao et al., 2015; Cvejoski et al., 2020). Their ability to encode fine-grained temporal structure, together with their capacity to reveal causal interactions between events in an interpretable manner, has made them indispensable not only in neuroscience and social media, but also in finance (Aït-Sahalia et al., 2015) and epidemiology (Chiang et al., 2022). Despite this centrality, the development of foundation models has followed a different trajectory. Large-scale pretraining first emerged in natural language processing, enabled by massive internet corpora, and has only recently been extended to dynamical systems, with recent work addressing ODEs (d’Ascoli et al., 2024; Mauel et al., 2025; 2026), MJPs (Berghaus et al., 2024), SDEs (Seifner et al., 2025a), and even specific applications in pharmacology (Marin et al., 2025). It is therefore ironic that event data — the very modality underlying the internet activity that made text-based pretraining possible — has not yet given rise to a corresponding foundation model for point processes. The present work takes a first step toward filling this gap by developing a foundation model explicitly designed for temporal point processes.
Marked Temporal Point Processes (MTPPs) (Daley and Vere-Jones, 2007; Rasmussen, 2018) are stochastic processes consisting of ordered occurrence times, each accompanied by a categorical mark specifying its type. Formally, the objective is to specify the conditional distribution of the next event time and mark given the history of the process up to the current time. The extensive point process literature explores different ways of encoding event histories and specifying the stochastic mechanism that governs new arrivals and their marks (Lin et al., 2024). A common approach is to represent this distribution through a conditional intensity function, which describes the instantaneous rate at which events of different types occur given the history. Traditionally, models such as the Hawkes process (Hawkes, 1971) define this conditional intensity as a superposition of self-exciting effects from past events. Forecasting is then carried out recursively using Ogata’s thinning algorithm, applied to the conditional intensity. Building on this cornerstone, more recent work has extended Hawkes processes by parameterizing event histories with neural architectures (Mei and Eisner, 2017), including recurrent neural networks (Du et al., 2016), attention mechanisms (Zhang et al., 2020), and Transformers (Zuo et al., 2021). These models are typically trained either via maximum likelihood — often requiring expensive integral evaluations — or through generative approaches that bypass intensity modeling altogether and directly sample future events conditional on the past (Kerrigan et al., 2024; Zeng et al., 2024). However, a fundamental limitation of these approaches is their lack of transferability: each new dataset requires retraining from scratch, forcing the model to relearn representations for each distinct dynamical regime.
In contrast, modern approaches to dynamical systems increasingly prioritize pretraining on synthetic data, yielding general models that can learn dynamics in-context. This paradigm offers a crucial advantage: practitioners no longer need to train models de novo for every dataset, but can instead obtain accurate characterizations in a zero-shot manner. Within this family, two variants have emerged: Prior-fitted Networks (PFNs) and Foundation Inference Models (FIMs). PFNs train networks to approximate predictive posterior distributions in a sequence-to-sequence or context-to-sequence fashion (Müller et al., 2022; Hollmann et al., 2022; Müller et al., 2025). FIMs, by contrast, focus on directly estimating the infinitesimal operators of dynamical system (e.g., drift and diffusion functions for SDEs), thereby retaining a degree of interpretability (Berghaus et al., 2024; Seifner et al., 2025a; b; Mauel et al., 2026). Access to these operators enables the explicit study of physically relevant observables such as entropy production, stationary distributions, and attractors.
In the context of point processes, we adapt the FIM pretraining paradigm to MTPPs in three steps. First, we define a broad family of conditional intensity functions, thereby inducing a diverse prior over MTPPs. This prior captures assumptions about excitatory and inhibitory effects between events, as well as the interaction structure across event types. Second, we sample MTPPs from this prior, generate synthetic event sequences, and construct tuples consisting of context event sequences, event histories, and their corresponding intensities, thereby creating a meta-learning task that amortizes inference across heterogeneous dynamics. Third, we train a neural network to recover conditional intensities from observed context. We summarize our contributions as follows:
  1. 1.
    Introduce a synthetic data generation framework for sampling event sequences from a broad prior distribution over MTPPs, with randomized base intensities, kernels, and interaction types (i.e., excitatory, inhibitory, neutral). We empirically demonstrate that this construction encodes a strong prior, enabling models trained on it to generalize across both in-distribution processes and real-world event data.
  2. 2.
    Train the first transformer-based recognition model capable of estimating in-context the conditional intensity functions of MTPPs, where history representations serve as queries and the encoded sequence context provides the keys and values.
  3. 3.
    Show that the resulting model achieves strong zero-shot performance across synthetic benchmarks and multiple real-world datasets, and that it can be rapidly finetuned on new event data.
2 Related Work
Here we provide a brief overview of temporal point processes. For detailed surveys and benchmarks of deep TPP models, including open challenges in history encoding, conditional intensity design, relational discovery, and learning strategies, see e.g., Lin et al. (2024) and  Xue et al. (2024).
While the mathematical theory of point processes is extensive (Daley and Vere-Jones, 2007; Kingman, 1992), work on temporal point processes (TPPs) in machine learning has crystallized around two central questions: (i) how should representations of past events be constructed, and (ii) how should the future be modeled (Lin et al., 2024). Early approaches, epitomized by the Hawkes process, addressed both questions using linear self-exciting kernels. A natural extension is the neural Hawkes process (Mei and Eisner, 2017), along with related recurrent formulations (Du et al., 2016), which rely on neural representations of past events, trained via likelihood maximization, and model future events using the thinning algorithm. Later work introduced more expressive architectures. Attention mechanisms (Zuo et al., 2021; Zhang et al., 2020; Yang et al., 2021) extend the memory horizon of TPPs, though at the cost of higher computational demand. Neural ODEs (Chen et al., 2018) have also been incorporated to better capture the irregular timing of events in latent representations (Song et al., 2024; Kidger et al., 2020).
To improve predictive accuracy over long horizons, different decompositions of the likelihood for future arrivals have been proposed (Rasmussen, 2018; Panos, 2024; Deshpande et al., 2021; Draxler et al., 2025). These works highlight the limitations of intensity-based inference, particularly when relying on thinning algorithms. Such limitations have motivated a shift toward generative models, which typically sample entire sequences. Approaches include optimal transport (Xiao et al., 2017), diffusion models (Zeng et al., 2024; Lüdke et al., 2023), and flow-matching methods (Kerrigan et al., 2024), often trading accuracy for interpretability. In contrast, traditional machine learning methods (Rasmussen, 2013; Malem-Shinitski et al., 2022) emphasize interpretability from the outset. A key advantage of the Hawkes process is that its excitation graph makes causal structure explicit (Xu et al., 2016; Wu et al., 2024), which has been especially relevant in neuroscience (Linderman and Adams, 2014; Truccolo et al., 2005) and in finance, where Hawkes models often serve as hidden drivers of observed activity (Aït-Sahalia et al., 2015). Applications extend more broadly, for instance to dynamics of text (Cvejoski et al., 2020; 2021), social online activity (Zhao et al., 2015) and operations research (Ojeda et al., 2021).
While Liu and Quan (2024) fine-tuned a pre-trained LLM for point-processes (with modest success), to the best of our knowledge, no prior work has presented a foundation model for point-processes that can operate zero-shot without further training.
3 Preliminaries
In this section, we recall the definition and basic properties of marked temporal point processes (Daley and Vere-Jones, 2007) and Hawkes processes (Hawkes, 1971; Laub et al., 2015). Additionally, we define the inference problem our proposed approach tackles.
Marked Temporal Point Processes: We consider marked temporal point processes (marked TPPs, or MTPPs) as simple point processes on $\mathbb{R}_{+}\times\mathcal{K}$, where $\mathcal{K}$ is a discrete and finite set of marks. The density $f$ of a sequence of events $\mathcal{S}=\{(t_{i},\kappa_{i})\}_{i=1}^{n}$ in the interval $[0,T]$, w.l.o.g. ordered by their time component $t_{i}\in\mathbb{R}_{+}$, factors into conditional densities
$f\left(\{(t_{i},\kappa_{i}\}_{i=1}^{n}\right)=\prod_{i=1}^{n}f\left((t_{i},\kappa_{i})\mid\mathcal{H}_{t_{i}}\right)=\prod_{i=1}^{n}f(t_{i}\mid\mathcal{H}_{t_{i}})f(\kappa_{i}\mid t_{i},\mathcal{H}_{t_{i}})\;,$ (1)
where $\mathcal{H}_{t}=\{(t_{i},\kappa_{i})\mid t_{i}<t\}\subset\mathcal{S}$ is the history strictly preceding $t$. By the last equality of equation 1, MTPPs may be characterized by dependent densities of the next-event time $f(t\mid\mathcal{H}_{t})$ and its event mark $f(\kappa\mid t,\mathcal{H}_{t})$. MTPPs are commonly represented by their piece-wise continuous conditional intensity function
$\lambda(t,\kappa\mid\mathcal{H}_{t})=\frac{f(t\mid\mathcal{H}_{t})}{1-\int_{t^{\prime}}^{t}f(s\mid\mathcal{H}_{s})ds}f(\kappa\mid t,\mathcal{H}_{t})=\lambda(t\mid\mathcal{H}_{t})f(\kappa\mid t,\mathcal{H}_{t}),$ (2)
where $t^{\prime}$ is the last event time in $\mathcal{H}_{t}$, or $t^{\prime}=0$ if $\mathcal{H}_{t}=\varnothing$. The conditional intensity function may be interpreted of the instantaneous rate of mark $\kappa$ occurring at $t$, conditioned of the history up to time $t_{i}$. Reversely, any such function $\lambda$, satisfying some mild conditions, defines the density of an MTPP on a set of events in an interval $[0,T]$ by
$f\left(\{(t_{i},\kappa_{i}\}_{i=1}^{n}\right)=\left[\prod_{i=1}^{n}\lambda(t_{i},\kappa_{i}\mid\mathcal{H}_{t_{i}})\right]\exp\left(-\int_{0}^{T}\lambda(s\mid\mathcal{H}_{s})ds\right).$ (3)
Collection of TPPs: A TPP is just an MTPP with a single mark. An MTPP can be viewed as a collection of TPPs per mark, interdependent through a joined history. Indeed, given an MTPP as above, the conditional intensity $\lambda_{\kappa}(t\mid\mathcal{H}_{t})=\lambda(t\mid\mathcal{H}_{t})f(\kappa\mid t,\mathcal{H}_{t})$ defines the marginal TPP for mark $\kappa\in\mathcal{K}$, that may depend on other marks via $\mathcal{H}_{t}$. Conversely, a collection of TPPs with conditional intensity $\lambda_{\kappa}$ per $\kappa\in\mathcal{K}$ can be joined to an MTPP. Using
$\lambda(t\mid\mathcal{H}_{t})=\sum_{\kappa\in\mathcal{K}}\lambda_{\kappa}(t\mid\mathcal{H}_{t})\quad\text{and}\quad f(\kappa\mid t,\mathcal{H}_{t})=\frac{\lambda_{\kappa}(t\mid\mathcal{H}_{t})}{\lambda(t\mid\mathcal{H}_{t})},$ (4)
in Eq. 2 defines the conditional intensity function $\lambda(t,\kappa\mid\mathcal{H}_{t})$ of an MTPP. In fact, $\lambda(t,\kappa\mid\mathcal{H}_{t})=\lambda_{\kappa}(t\mid\mathcal{H}_{t})$. In contrast to some other neural methods (Du et al., 2016), which estimate $\lambda(t\mid\mathcal{H}_{t})$ and $f(\kappa\mid t,\mathcal{H}_{t})$, we design our model to parametrize TPPs per mark, conditioned on the joined history of all marks.
Hawkes Processes: A Hawkes MTPP with marks $\mathcal{K}$ is defined by the conditional intensity
$\lambda(t,\kappa\mid\mathcal{H}_{t})=\max\left(0,\mu_{\kappa}(t)+\sum_{(t^{\prime},\kappa^{\prime})\in\mathcal{H}_{t}}\gamma_{\kappa\kappa^{\prime}}(t-t^{\prime})\right),$ (5)
where $\{\mu_{k}\}_{\kappa\in\mathcal{K}}$ is a set of time-dependent base intensity functions, and $\{\gamma_{\kappa\kappa^{\prime}}\}_{\kappa,\kappa^{\prime}\in\mathcal{K}}$ is a set of interaction kernels, specifying the influence of mark $\kappa^{\prime}$ on mark $\kappa$. If $\gamma_{\kappa\kappa^{\prime}}$ is positive, the influence of $\kappa^{\prime}$ on $\kappa$ is called excitatory or exciting, otherwise it is called inhibitory or limiting.
Simulation: We use Ogata’s modified thinning algorithm (Ogata, 1981) to generate synthetic training data from MTPP, and to simulate processes inferred by our model.
Inference Problem: Let $\mathcal{C}=\{\mathcal{S}^{j}\}_{j=1}^{m}$ be a collection of $m$ event sequences $\mathcal{S}^{j}=\{(t_{i}^{j},\kappa_{i}^{j})\}_{i=1}^{n_{j}}$ observed from some system. Our objective is to predict or simulate the next event and estimate the likelihood of a (previously unseen) sequence $\mathcal{S}$, assuming an MTPP model. Previous neural methods train an autoregressive encoding network on $\mathcal{C}$ that compresses the history $\mathcal{H}_{t}$ of $\mathcal{S}$ into some embedding $\mathbf{h}_{t}$ for a neural estimate $\hat{\lambda}(t,\kappa\mid\mathbf{h}_{t})$ of the conditional intensity. In contrast, we pretrain a deep neural network model to estimate in-context $\hat{\lambda}$, given a history of events, from a collection of context event sequences. Once trained, the model can be applied to any $\mathcal{C}$ and $\mathcal{H}_{t}$, without any further training.
4 Foundation Inference Models for Point Processes
In this section, we present a novel in-context learning method for the MTPP intensity inference problem. In a two-step approach, we first generate a large set of marked event sequences from parametrized MTPPs, sampled from a broad distribution over MTPPs. This yields train data for a neural network recognition model, trained to estimate the underlying known, ground-truth intensity functions. Such pretrained inference model can be applied directly to real-world problems, or swiftly finetuned for improved performance.
4.1 Synthetic Dataset Generation
To construct a synthetic dataset of MTPPs, we define a distribution over MTPPs via a distribution over conditional intensity functions of the form
$\lambda(t,\kappa\mid\mathcal{H}_{t})=\max\left(0,\mu_{\kappa}(t)+\sum_{(t^{\prime},\kappa^{\prime})\in\mathcal{H}_{t}}z_{\kappa\kappa^{\prime}}\gamma_{\kappa\kappa^{\prime}}(t-t^{\prime})\right)\;.$ (6)
Given a sample from this distribution, we simulate a large set of marked event sequences and record the corresponding conditional intensities for training.
We use Eq. 6 to parameterize several classes of processes:
  • The classical Hawkes process, for which $\mu_{k}$ is constant in time and $\gamma_{\kappa\kappa^{\prime}}(t)$ is an exponential function;
  • The Poisson process, for which $\mu_{k}$ is constant in time and $\gamma_{\kappa\kappa^{\prime}}=0$;
  • Periodic processes, for which $\mu_{k}(t)$ is a positive sinusoidal function;
  • Processes with high initial excitation, for which we choose $\mu_{k}(t)$ to follow a Gamma distribution;
  • Processes with non-monotonic, shifted kernels, for which we choose $\gamma_{\kappa\kappa^{\prime}}(t)$ to follow a Rayleigh distribution.
A complete overview of the process families and their hyperparameters is provided in Table 6 in the Appendix. Furthermore, Figure 7 in the Appendix reports summary statistics of our simulated pretraining distribution, and confirms that it is broad enough to cover the real-world datasets depicted in Figures 8 to 12, also in the Appendix.
For each process, we randomly sample $z_{\kappa\kappa^{\prime}}\in\{-1,0,1\}$ for ${\kappa,\kappa^{\prime}}\in\mathcal{K}$ to cover excitatory ($z_{\kappa\kappa^{\prime}}=1$), inhibitory ($z_{\kappa\kappa^{\prime}}=-1$) and non-influencing ($z_{\kappa\kappa^{\prime}}=0$) interactions.
4.2 Foundation Inference Model Architecture
Figure 1: Schematic representation of FIM-PP. A context of marked event sequences $\mathcal{S}^{j}$ is encoded by a self-attentive transformer encoder. The result is further processed by a transformer decoder, using a history $\mathcal{H}_{t}$ of marked events before time $t$ as queries. The final embedding is joined with an encoding of mark $\kappa$. The results is projected to a set of parameters that determine the value of the conditional intensity function $\hat{\lambda}$ evaluated at $(t,\kappa)$.
We now present the architecture of FIM-PP, a pretrained deep neural network for inference of MTPPs from sets $\mathcal{C}=\{\mathcal{S}^{j}\}_{j=1}^{m}$ of marked event sequences $\mathcal{S}^{j}=\{(t^{j}_{i},\kappa^{j}_{i})\}_{i=1}^{n_{j}}$. FIM-PP processes the context sequences $\mathcal{C}$ and estimates the conditional intensity function $\hat{\lambda}$ of an MTPP that describes the observed dynamics. Following previous intensity-based methods (Zhang et al., 2020; Zuo et al., 2021), $\hat{\lambda}$ is implemented by a flexible parametrized function family $\hat{\lambda}(\cdot,\kappa\mid\mathcal{H}_{\cdot})$ for all $\kappa\in\mathcal{K}$. FIM-PP estimates its parameters by encoding the history $\mathcal{H}_{t}=\{(t^{\text{{hist}}}_{i},\kappa^{\text{{hist}}}_{i})\}_{i=1}^{n_{\text{{hist}}}}$ before time $t>t^{\text{{hist}}}_{n_{\text{{hist}}}}$, subject to the processed context sequences. Figure 1 depicts a schematic representation of this approach.
To cover applications in different time scales, FIM-PP instance normalizes its inputs and renormalizes $\hat{\lambda}$ accordingly. Appendix C provides the details. Once trained, FIM-PP can be applied for all counts of marks $\lvert\mathcal{K}\rvert$ up to some fixed upper bound, similar to in-context methods in other domains (d’Ascoli et al., 2024; Berghaus et al., 2024).
We denote linear projections by $\phi$, feed-forward neural networks by $\Phi$, attention layers with residual connections by $\psi$, transformer encoders by $\Psi_{\text{{enc}}}$ and decoders by $\Psi_{\text{{dec}}}$. Let $E\in\mathbb{N}$ denote the model’s embedding dimension.
Context Encoding: To encode $\mathcal{C}$, we combine encodings of individual sequences $\mathcal{S}^{j}$. Recognizing the importance of inter-observation times for the inference problem, we consider $\Delta t_{i}^{j}=t_{i}^{j}-t_{i-1}^{j}$ as an additional feature, identifying $t_{0}^{j}=0$. To encode $\mathcal{S}^{j}$, we first embed the features $(t_{i}^{j},\kappa_{i}^{j},\Delta t_{i}^{j})$ of the $i$-th event in sequence $j$ into embeddings
$\mathbf{u}_{i}^{j}=\phi_{t}(t_{i}^{j})+\phi_{\kappa}(\kappa_{i}^{j})+\phi_{\Delta t}(\Delta t_{i}^{j})\,\in\,\mathbb{R}^{E}\;.$ (7)
Sinusoidal output activations from Shukla and Marlin (2020) enhance the networks $\phi_{t}$ and $\phi_{\Delta t}$. Let $\mathbf{S}^{j}=[\mathbf{u}_{1}^{j},\dots,\mathbf{u}_{n_{j}}^{j}]\in\mathbb{R}^{n_{j}\times E}$ denote the matrix of embedding of sequence $\mathcal{S}^{j}$. We extract a context sequence embedding $\mathbf{c}_{j}\in\mathbb{R}^{E}$ by applying a transformer encoder $\tilde{\mathbf{S}}^{j}=\Psi_{\text{{enc}}}^{\text{{cont}}}(\mathbf{S}^{j})\in\mathbb{R}^{n_{j}\times E}$, followed by fixed-query attention
$\mathbf{c}_{j}=\psi^{\text{{cont}}}(\mathbf{q}^{\text{{cont}}},\tilde{\mathbf{S}}^{j},\tilde{\mathbf{S}}^{j})\,\in\mathbb{R}^{E}\,$, (8)
where $\mathbf{q}^{\text{{cont}}}\in\mathbb{R}^{E}$ is a learnable query.
We emphasize that we use an hierarchical approach where every sequence gets processed independently by $\Psi_{\text{{enc}}}^{\text{{cont}}}$ and encoded into a single embedding $\mathbf{c}_{j}$. This is much more scalable compared to combining all events into a single long sequence. The embeddings of all sequences $\mathbf{C}=[\mathbf{c}_{1},\dots,\mathbf{c}_{m}]\in\mathbb{R}^{m\times E}$ are finally combined by another transformer encoder $\tilde{\mathbf{C}}=\Psi_{\text{{enc}}}^{\text{{comb}}}(\mathbf{C})\in\mathbb{R}^{m\times E}$.
Context-aware History Encoding: To encode the history $\mathcal{H}_{t}$ of events prior to time $t>t^{\text{{hist}}}_{n_{\text{{hist}}}}$, we embed each tuple $(t^{\text{{hist}}}_{i},\kappa^{\text{{hist}}}_{i},\Delta t_{i}^{\text{{hist}}})$ into feature vectors $\mathbf{H}=[\mathbf{u}_{1}^{\text{{hist}}},\dots,\mathbf{u}_{n_{\text{{hist}}}}^{\text{{hist}}}]\in\mathbb{R}^{n_{\text{{hist}}}\times E}$, reusing the networks from Eq. 7. These history embeddings serve as the queries of a transformer decoder $\Psi_{\text{{dec}}}^{\text{{hist}}}$, which attends to the context representation $\tilde{\mathbf{C}}$ (used as keys and values), yielding a unified encoding $\mathbf{h}^{\text{{hist}}}_{t}$ that integrates both history and context:
$\mathbf{h}^{\text{{hist}}}_{t}=\Psi_{\text{{dec}}}^{\text{{hist}}}(\mathbf{H},\tilde{\mathbf{C}})\in\mathbb{R}^{E}.$ (9)
Intensity Parametrization: To extract intensity functions for all marks from $\mathbf{h}^{\text{{hist}}}_{t}$, we concatenate $\mathbf{h}^{\text{{hist}}}_{t}$ with a (linear) encoding of $\kappa^{\prime}$ to form $\mathbf{v}^{\mathcal{H}_{t}}_{\kappa^{\prime}}=[\mathbf{h}^{\text{{hist}}}_{t},\phi_{\kappa}^{\text{{hist}}}(\kappa^{\prime})]\in\mathbb{R}^{2E}$, and project it to non-negative parameter estimates
$\hat{\alpha}_{\kappa^{\prime}}^{\mathcal{H}_{t}}=\Phi_{\alpha}(\mathbf{v}^{\mathcal{H}_{t}}_{\kappa^{\prime}})\in\mathbb{R}_{+},\quad\hat{\beta}_{\kappa^{\prime}}^{\mathcal{H}_{t}}=\Phi_{\beta}(\mathbf{v}^{\mathcal{H}_{t}}_{\kappa^{\prime}})\in\mathbb{R}_{+}\quad\text{and}\quad\hat{\mu}_{\kappa^{\prime}}^{\mathcal{H}_{t}}=\Phi_{\mu}(\mathbf{v}^{\mathcal{H}_{t}}_{\kappa^{\prime}})\in\mathbb{R}_{+}.$ (10)
We enforce non-negativity via softplus output activations. These parameters define our neural conditional intensity estimate: [^2] Note that the functional form of $\hat{\lambda}$ is similar to the conditional intensity in Zhang et al. (2020).
$\hat{\lambda}(t,\kappa^{\prime}\mid\mathcal{H}_{t})=\hat{\mu}_{\kappa^{\prime}}^{\mathcal{H}_{t}}+(\hat{\alpha}_{\kappa^{\prime}}^{\mathcal{H}_{t}}-\hat{\mu}_{\kappa^{\prime}}^{\mathcal{H}_{t}})\exp\left(-\hat{\beta}_{\kappa^{\prime}}^{\mathcal{H}_{t}}(t-t_{n_{\text{{hist}}}})\right).$ (11)
This parametrization is both flexible and interpretable. Indeed, immediately after incorporating a new event into the history, the intensity jumps to $\hat{\alpha}_{\kappa^{\prime}}^{\mathcal{H}_{t}}$. Over long inter-event intervals, the intensity relaxes toward $\hat{\mu}_{\kappa^{\prime}}^{\mathcal{H}_{t}}$. The relaxation rate is governed by $\hat{\beta}_{\kappa^{\prime}}^{\mathcal{H}_{t}}$.
Although this parametrization may appear tailored to Hawkes-type dynamics, its parameters $\hat{\mu}_{\kappa^{\prime}}^{\mathcal{H}_{t}}$, $\hat{\alpha}_{\kappa^{\prime}}^{\mathcal{H}_{t}}$, and $\hat{\beta}_{\kappa^{\prime}}^{\mathcal{H}_{t}}$ are themselves history- and mark-dependent. As a result, the model can represent rich local intensity behaviors, including localized triggering patterns (e.g., Rayleigh- or power-law-like kernels) and time-dependent baseline intensities. We highlight this empirically in Section 5.
Figure 2: Example intensity estimates of FIM-PP on a synthetic Hawkes process with three marks, constant base intensity and exponential decaying kernels (left) and a real-world Retweet dataset (right). Each row contains the intensity for one mark. Events of the same mark are colored magenta, while events for other marks are gray. For the Hawkes process, the model (blue line) estimate matches the ground-truth intensity level (black dashed line) closely. For the Retweet data, FIM-PP estimates a mixture of many excitatory and a few inhibitory interactions.
Training: To train FIM-PP on a set of sequences $\mathcal{C}_{\lambda}$ from our synthetic pretraining data, we select a target sequence $\mathcal{T}\in\mathcal{C}_{\lambda}$ to provide a history of events and use remaining sequences $\mathcal{C}_{\lambda}\setminus\{\mathcal{T}\}$ as context. We subsample $\mathcal{C}_{\lambda}$, truncate sequences and vary the number of marks throughout training, which enables us to apply a pretrained FIM-PP in a wide range of (real-world) settings. Our train objective is the next-event negative log likelihood of the target sequence:
$\mathcal{L}_{\text{NLL}}=\sum_{\kappa\in\mathcal{K}}\int_{0}^{T}\hat{\lambda}(s,\kappa\mid\mathcal{H}_{s})ds-\sum_{(t,\kappa)\in\mathcal{T}}\hat{\lambda}(t,\kappa\mid\mathcal{H}_{t}).$ (12)
We remark that we also experimented with a supervised sMAPE loss, the results of which can be found in the Appendix Tables, starting at Table 3. We found that this approach performs similarly well but is more computationally expensive and less flexible, which is why we use the NLL model in the main text. Appendix D discusses the training of FIM-PP in greater detail.
Finetuning: FIM-PP can be finetuned on the train split $\mathcal{C}$ of an evaluation dataset, minimizing $\mathcal{L}_{\text{NLL}}$. For each iteration, a random sequence $\mathcal{T}\in\mathcal{C}$ in the train split is selected as the target sequence. The remaining sequences $\mathcal{C}\setminus\{\mathcal{T}\}$ serve as context. Finetuning progress is monitored by processing target sequences from the validation split, given the train split context.
5 Experiments
In this section, we repeat the experiments by Zeng et al. (2024), who introduced CDiff, a recent state-of-the-art diffusion-based marked event sequence forecasting model. They compare their method against a range of intensity-based and intensity-free baselines, on common benchmark datasets, evaluated on standard metrics. Moreover, they made their code and exact evaluation dataset splits available [^3] https://github.com/networkslab/cdiff, which allows us to directly compare against their results. In the following, we recall their experimental setup, describe the pretraining and application of FIM-PP, before presenting and analyzing our results.
5.1 Experimental Setup
Prediction Task: Given a sequence of events $\mathcal{S}=\{(t_{i},\kappa_{i})\}_{i=1}^{n}$, the task is to predict the next $N\in\mathbb{N}$ events following $\mathcal{S}$, where $N$ is the prediction horizon length. We denote the ground-truth continuation by $\mathcal{S}_{\text{{final}}}^{N}$ and the model prediction by $\hat{\mathcal{S}}_{\text{{final}}}^{N}$. We refer to the case $N=1$ as next-event prediction and to $N>1$ as multi-event prediction.
[…]
Evaluation Data: We benchmark on five widely used real-world datasets: Taxi, Taobao, StackOverflow, Amazon, and Retweet. These datasets vary in the number of marks, sequence lengths, and event counts, making them a strong testbed for evaluating the broad applicability of FIM-PP. We use the preprocessing and train/test/validation splits of Zeng et al. (2024). Appendix E contains further details, including dataset statistics and original sources.
Baselines: We compare FIM-PP against methods falling into two categories: models that learn joint distributions over multiple events, and autoregressive approaches such as FIM-PP. The first category includes Dual-TPP (Deshpande et al., 2021), HYPRO (Xue et al., 2022), and the Cross-diffusion Model (CDiff) (Zeng et al., 2024). The second category further splits into intensity-based and intensity-free approaches. Intensity-based baselines are the Neural Hawkes Process (NHP) (Mei and Eisner, 2017), and the Attentive Neural Hawkes Process (A-NHP) (Mei et al., 2022). Intensity-free baselines are the Intensity-Free Temporal Point Process (IFTPP) (Shchur et al., 2020), and the Temporal Conditional Diffusion Denoising Model (TCDDM) (Lin et al., 2022).
(a) Mean performance metrics across five different datasets for different horizon lengths.
(b) Next event prediction on Taxi after fine-tuning
Figure 3: (a) shows that FIM-PP (zs) is competitive but slightly worse than the baseline models. FIM-PP (f) however performs best among all horizon lengths. (b) shows that FIM-PP (f) also reliably captures patterns in the Taxi dataset.
5.2 Pretraining, Fine-tuning and Intensity Predictions
Pretraining. We pretrain a single FIM-PP on a synthetic dataset containing 14.414.4M events, simulated from 7272K point processes of diverse kernels, and sparsity levels, and varying number of marks, sequences, and events. Appendix B contains the details. The model has 1616M parameters and supports up to $\lvert\mathcal{K}\rvert=22$ marks to cover all evaluation datasets. Further pretraining details are provided in Appendix D.
Finetuning. In zero-shot mode, we apply the pretrained model directly to all evaluation datasets, and label the results by FIM-PP (zs). We also experiment with finetuning FIM-PP on the train split of all evaluation datasets, and label these results by FIM-PP (f). FIM-PP utilizes up to $2000$ sequences from the train split of an evaluation dataset as context, limited only by the maximum number of sequences seen during training. The ablations in Figure 6 also indicate that typically much less than 2000 context paths would be sufficient. The sequences in the test split are given as context to FIM-PP. Finally, for multi-event prediction, FIM-PP simulates events autoregressively, similar to other (intensity-based) baselines.
Due to the relatively small parameter size of FIM-PP and the pre-trained prior, fine-tuning can be performed quickly. For all datasets, fine-tuning was achieved in a few minutes, requiring not more than 11 GB of memory. This means that fine-tuning FIM-PP does not take longer than training the baseline models from scratch, which reportedly takes up to 4 hours (Xue et al., 2024, Sec. D.1). Fine-tuning is therefore feasible for users with small computational resources. Further speed-ups might be possible with LoRA (Hu et al., 2022).
Intensity Predictions. Figure 2 shows intensities inferred by FIM-PP in zero-shot mode, both from an unseen synthetic Hawkes process (in-distribution generalization), and from the Retweet dataset (out-of-distribution generalization). Moreover, Figure 4 in the Appendix illustrates the performance of FIM-PP (zs) on various synthetic datasets and emphasizes that FIM-PP (zs) also generalizes well for processes with powerlaw kernels, a kernel type that was not present in the pretraining dataset. This confirms that our parameterization is general enough. In what follows, we quantitatively evaluate FIM-PP.
Table 1: Performance on four real-world datasets, predicting $N=20$ events. Results for baseline methods were extracted from Zeng et al. (2024). We report mean and standard deviation over $10$ trials for two metrics. Best results are bold.
Taxi StackOverflow Amazon Retweet
Method $\Delta t OTD sMAPE$$\Delta t$ $\Delta t OTD sMAPE$$\Delta t$
HYPRO $21.60{\text{{$\pm 0.20$}}}$ $93.8{\text{{$\pm 0.4$}}}$ $42.40{\text{{$\pm 0.20$}}}$ $111.00{\text{{$\pm 0.60$}}}$ $38.6{\text{{$\pm 0.5$}}}$ $82.5{\text{{$\pm 0.8$}}}$ $61.03{\text{{$\pm 0.09$}}}$ $106.11{\text{{$\pm 1.51$}}}$
Dual-TPP $24.48{\text{{$\pm 0.38$}}}$ $95.2{\text{{$\pm 0.2$}}}$ $41.75{\text{{$\pm 0.20$}}}$ $117.58{\text{{$\pm 0.42$}}}$ $42.6{\text{{$\pm 0.7$}}}$ $86.5{\text{{$\pm 2.0$}}}$ $61.10{\text{{$\pm 0.10$}}}$ $106.90{\text{{$\pm 1.29$}}}$
A-NHP $24.76{\text{{$\pm 0.22$}}}$ $97.4{\text{{$\pm 0.4$}}}$ $42.59{\text{{$\pm 0.41$}}}$ $108.54{\text{{$\pm 0.53$}}}$ $39.5{\text{{$\pm 0.3$}}}$ $84.3{\text{{$\pm 1.8$}}}$ $60.63{\text{{$\pm 0.10$}}}$ $107.23{\text{{$\pm 1.29$}}}$
NHP $25.11{\text{{$\pm 0.27$}}}$ $96.5{\text{{$\pm 0.5$}}}$ $43.79{\text{{$\pm 0.15$}}}$ $116.95{\text{{$\pm 0.40$}}}$ $42.6{\text{{$\pm 0.3$}}}$ $92.1{\text{{$\pm 1.6$}}}$ $60.95{\text{{$\pm 0.08$}}}$ $107.08{\text{{$\pm 1.40$}}}$
IFTPP $24.05{\text{{$\pm 0.61$}}}$ $95.7{\text{{$\pm 0.8$}}}$ $46.28{\text{{$\pm 0.89$}}}$ $115.12{\text{{$\pm 0.63$}}}$ $43.8{\text{{$\pm 0.2$}}}$ $90.9{\text{{$\pm 1.6$}}}$ $61.72{\text{{$\pm 0.15$}}}$ $106.71{\text{{$\pm 1.62$}}}$
TCDDM $22.15{\text{{$\pm 0.53$}}}$ $90.6{\text{{$\pm 0.6$}}}$ $42.13{\text{{$\pm 0.59$}}}$ $107.66{\text{{$\pm 0.93$}}}$ $42.2{\text{{$\pm 0.2$}}}$ $83.8{\text{{$\pm 1.5$}}}$ $60.50{\text{{$\pm 0.09$}}}$ $106.05{\text{{$\pm 0.61$}}}$
CDiff $21.01{\text{{$\pm 0.16$}}}$ $88.0{\text{{$\pm 0.2$}}}$ $41.25{\text{{$\pm 1.40$}}}$ $106.18{\text{{$\pm 0.34$}}}$ $37.7{\text{{$\pm 0.2$}}}$ $82.0{\text{{$\pm 1.9$}}}$ $60.66{\text{{$\pm 0.10$}}}$ $106.18{\text{{$\pm 1.12$}}}$
FIM-PP (zs) $23.15{\text{{$\pm 0.07$}}}$ $\mathbf{76.8}{\text{{$\pm 0.4$}}}$ $49.26{\text{{$\pm 0.06$}}}$ $96.36{\text{{$\pm 0.05$}}}$ $46.2{\text{{$\pm 0.1$}}}$ $128.6{\text{{$\pm 0.4$}}}$ $60.24{\text{{$\pm 0.16$}}}$ $99.07{\text{{$\pm 0.39$}}}$
FIM-PP (f) $\mathbf{17.91}{\text{{$\pm 0.12$}}}$ $76.8{\text{{$\pm 0.5$}}}$ $\mathbf{39.80}{\text{{$\pm 0.04$}}}$ $\mathbf{88.25}{\text{{$\pm 0.19$}}}$ $\mathbf{37.2}{\text{{$\pm 0.1$}}}$ $\mathbf{81.2}{\text{{$\pm 0.1$}}}$ $\mathbf{59.44}{\text{{$\pm 0.08$}}}$ $\mathbf{87.59}{\text{{$\pm 0.17$}}}$
5.3 Multi-Event Prediction
[…]
Figure 3(a) summarizes performance across horizon lengths by averaging results over all datasets. In aggregate, FIM-PP (zs) performs on par with the baselines, whereas FIM-PP (f) consistently outperforms them, in agreement with our previous analysis. We attribute the effectiveness of finetuning to two factors: (i) the strong prior encoded in the model weights through pretraining on our synthetic distribution, which provides a favorable initialization for finetuning; and (ii) the flexibility of the foundation model architecture, which enables direct access to patterns in the training split at evaluation time, since these patterns can be retrieved from the context. The first point is supported by Figure 5 in the Appendix, which shows that the pretrained model converges faster and achieves better final performance than the same architecture trained from scratch.
5.4 Next-Event Prediction
The next-event prediction task is a special case of multi-event prediction, but it differs in nature. Whereas multi-event prediction requires estimating the distribution over a set of future events, next-event prediction focuses on accurately forecasting a single event. This distinction is also reflected in the evaluation metrics (see Appendix F) [^4] For instance, mark accuracy (Acc) targets the correctness of one event, while RMSEee compares histograms over multiple events..
Table 2 reports next-event prediction results ($N=1$) on two real-world datasets. FIM-PP (zs) performs well on event-time prediction for Taxi, but struggles with mark accuracy on Taxi and with both event-time and mark accuracy on Taobao. These difficulties can be explained by dataset-specific patterns: sequences in Taxi often alternate consistently between two marks, a pattern unlikely to appear in our point process prior. In contrast, Taobao is heavily dominated by a single mark and occasionally exhibits long waiting times. Again, patterns not covered by our pretraining distribution.
Compared to the baselines, which easily incorporate such patterns during training, FIM-PP (zs) cannot reproduce them accurately from the context alone. Appendix G further analyzes these out-of-distribution patterns.
When finetuned, however, FIM-PP can adapt to (some) of these characteristics. Table 2 shows substantial improvements in mark prediction accuracy (Acc) after finetuning, and Figure 3 illustrates that FIM-PP (f) successfully recovers the alternating pattern in the Taxi dataset. Nevertheless, FIM-PP (f) still lags behind the baselines in mark prediction accuracy (Acc). In Appendix G, we suggest broadening the synthetic pretraining distribution to better capture such distinctive patterns upfront.
Table 2: Next-event prediction performance on two real-world datasets, displayed with mean and standard deviation over $10$ trials. Results for baseline methods were extracted from Zeng et al. (2024) (we remark that they did not report next-event prediction results for the other datasets). Best results are bold.
Taxi Taobao
Method $Delta t RMSEΔt\Delta t Acc sMAPE$$\Delta t$
A-NHP $0.32$ $\pm 0.00$ $\mathbf{0.91}$ $\pm 0.01$ $85.13$ $\pm 0.26$ $0.53$ $\pm 0.00$ $0.47$ $\pm 0.01$ $129.13$ $\pm 1.35$
Dual-TPP $0.34$ $\pm 0.01$ $\mathbf{0.91}$ $\pm 0.01$ $89.12$ $\pm 0.75$ $0.53$ $\pm 0.01$ $0.47$ $\pm 0.02$ $131.43$ $\pm 1.99$
NHP $0.34$ $\pm 0.01$ $\mathbf{0.91}$ $\pm 0.01$ $90.63$ $\pm 0.61$ $0.53$ $\pm 0.00$ $0.46$ $\pm 0.01$ $133.69$ $\pm 2.25$
IFTPP $0.38$ $\pm 0.01$ $0.90$ $\pm 0.01$ $90.03$ $\pm 0.47$ $0.53$ $\pm 0.01$ $0.45$ $\pm 0.01$ $\mathbf{126.01}$ $\pm 1.48$
CDiff $0.34$ $\pm 0.01$ $\mathbf{0.91}$ $\pm 0.00$ $87.12$ $\pm 0.61$ $\mathbf{0.52}$ $\pm 0.01$ $\mathbf{0.48}$ $\pm 0.00$ $127.12$ $\pm 1.36$
FIM-PP (zs) $\mathbf{0.15}$ $\pm 0.00$ $0.41$ $\pm 0.03$ $69.37$ $\pm 0.91$ $1.41$ $\pm 0.03$ $0.09$ $\pm 0.01$ $163.34$ $\pm 0.51$
FIM-PP (f) $\mathbf{0.15}$ $\pm 0.00$ $0.69$ $\pm 0.01$ $\mathbf{63.02}$ $\pm 0.48$ $9.31$ $\pm 0.15$ $0.39$ $\pm 0.01$ $138.46$ $\pm 2.69$
6 Conclusions
In this work, we introduced FIM-PP, the first Foundation Inference Model for inferring marked temporal point processes (MTPPs) from real-world data. Our experiments show that a single FIM-PP, pretrained only on synthetic MTPP data, can already match the predictive performance of existing intensity-based MTPP methods in zero-shot mode, i.e., without any further training. Finetuning on target data further improves results within only a few iterations, enabling FIM-PP to outperform competing methods on the majority of evaluated tasks.
Limitations: Although our pretraining distribution is broad, it is not universal and therefore cannot capture many real-world patterns. As a result, zero-shot performance may degrade under distribution shift. FIM-PP is further constrained by the maximum number of marks $\lvert\mathcal{K}\rvert$ and the maximum sequence length used during training. When these limits are exceeded, the model may not fully exploit the available context.
Future Work: An important direction for future work is to broaden the pretraining distribution beyond the parametrization in Eq. 6, with the goal of capturing a wider range of data patterns in zero-shot mode, and providing an even stronger initialization for finetuning. In addition, intensity-free methods have recently shown strong predictive performance (Panos, 2024). We plan to investigate how such methods can be incorporated into our amortized in-context learning framework.
7 Reproducibility Statement
Our core methodology consists of two parts: synthetically generated training data and a foundation inference model for marked temporal point process. Data generation is described extensively in Section 4.1 and complemented by Appendix B, which covers the exact hyperparameters and design choices required to reproduce our training dataset. Section 4.2 describes the architecture of FIM-PP. Training details, including hyperparameter choices and submodule sizes, are described in Appendix D. Our pretrained model, repository, and tutorials are available online [^5] https://fim4science.github.io/OpenFIM/intro.html .
The real-world datasets used in our experiments are described in Appendix E, including dataset sizes and numbers of marks. For data sourcing and preprocessing, we follow Zeng et al. (2024), as discussed in Appendix E. Finally, the evaluation metrics used in all experiments are described in Appendix F.
Acknowledgments
This research has been funded by the Federal Ministry of Education and Research of Germany and the state of North-Rhine Westphalia as part of the Lamarr Institute for Machine Learning and Artificial Intelligence. Additionally, César Ojeda was supported by Deutsche Forschungsgemeinschaft (DFG) – Project-ID 318763901 – SFB1294.
Appendix A Additional Results
A.1 Performance on various synthetic Datasets
In this section we highlight the performance of FIM-PP (zs) on various synthetic datasets coming from different processes. Figure 4 compares the estimated intensity to the ground-truth on some synthetic processes, including the powerlaw kernel. The model captures the essence of all processes, including the out-of-distribution powerlaw kernel.
Figure 4: Intensity predictions of FIM-PP (zs) on synthetic datasets of four different process types. We remark that the model has not been trained on powerlaw kernels but still predicts them with decent accuracy.
A.2 Long-Horizon Prediction
[…]
Table 3: Prediction of $N=20$ events in test sequences of five real-world datasets. Error-bars indicate the standard deviation over $10$ trials. Results for the baseline methods were extracted from Zeng et al. (2024). Best results are bold.
[…]
Taxi HYPRO $21.653{\text{{$\pm 0.163$}}}$ $1.231{\text{{$\pm 0.015$}}}$ $0.372{\text{{$\pm 0.004$}}}$ $93.803{\text{{$\pm 0.454$}}}$
Dual-TPP $24.483{\text{{$\pm 0.383$}}}$ $1.353{\text{{$\pm 0.037$}}}$ $0.402{\text{{$\pm 0.006$}}}$ $95.211{\text{{$\pm 0.187$}}}$
A-NHP $24.762{\text{{$\pm 0.217$}}}$ $1.276{\text{{$\pm 0.015$}}}$ $0.430{\text{{$\pm 0.003$}}}$ $97.388{\text{{$\pm 0.381$}}}$
NHP $25.114{\text{{$\pm 0.268$}}}$ $1.297{\text{{$\pm 0.019$}}}$ $0.399{\text{{$\pm 0.040$}}}$ $96.459{\text{{$\pm 0.521$}}}$
IFTPP $24.053{\text{{$\pm 0.609$}}}$ $1.364{\text{{$\pm 0.032$}}}$ $0.384{\text{{$\pm 0.005$}}}$ $95.719{\text{{$\pm 0.779$}}}$
TCDDM $22.148{\text{{$\pm 0.529$}}}$ $1.309{\text{{$\pm 0.030$}}}$ $0.382{\text{{$\pm 0.019$}}}$ $90.596{\text{{$\pm 0.574$}}}$
CDiff $21.013{\text{{$\pm 0.158$}}}$ $1.131{\text{{$\pm 0.017$}}}$ $0.351{\text{{$\pm 0.004$}}}$ $87.993{\text{{$\pm 0.178$}}}$
FIM-PP (zs) (NLL) $23.145$ $\pm 0.073$ $1.421$ $\pm 0.014$ $\mathbf{0.277}$ $\pm 0.000$ $\mathbf{76.765}$ $\pm 0.386$
FIM-PP (zs) (S-MAPE) $41.543$ $\pm 0.026$ $3.986$ $\pm 0.009$ $0.297$ $\pm 0.000$ $90.149$ $\pm 0.366$
FIM-PP (f) (NLL) $17.914$ $\pm 0.117$ $\mathbf{0.705}$ $\pm 0.006$ $0.314$ $\pm 0.004$ $\mathbf{76.828}$ $\pm 0.549$
FIM-PP (f) (S-MAPE) $\mathbf{16.493}$ $\pm 0.034$ $\mathbf{0.707}$ $\pm 0.004$ $\mathbf{0.278}$ $\pm 0.001$ $\mathbf{76.538}$ $\pm 0.214$
Taobao HYPRO $\mathbf{44.336}{\text{{$\pm 0.127$}}}$ $2.710{\text{{$\pm 0.021$}}}$ $0.594{\text{{$\pm 0.030$}}}$ $134.922{\text{{$\pm 0.473$}}}$
Dual-TPP $47.324{\text{{$\pm 0.541$}}}$ $3.237{\text{{$\pm 0.049$}}}$ $0.871{\text{{$\pm 0.005$}}}$ $141.687{\text{{$\pm 0.431$}}}$
A-NHP $45.555{\text{{$\pm 0.345$}}}$ $2.737{\text{{$\pm 0.021$}}}$ $0.708{\text{{$\pm 0.010$}}}$ $134.582{\text{{$\pm 0.920$}}}$
NHP $48.131{\text{{$\pm 0.297$}}}$ $3.355{\text{{$\pm 0.030$}}}$ $0.837{\text{{$\pm 0.009$}}}$ $137.644{\text{{$\pm 0.764$}}}$
IFTPP $45.757{\text{{$\pm 0.287$}}}$ $3.193{\text{{$\pm 0.043$}}}$ $0.575{\text{{$\pm 0.012$}}}$ $127.436{\text{{$\pm 0.606$}}}$
TCDDM $45.563{\text{{$\pm 0.889$}}}$ $2.850{\text{{$\pm 0.058$}}}$ $0.569{\text{{$\pm 0.015$}}}$ $126.512{\text{{$\pm 0.491$}}}$
CDiff $44.621{\text{{$\pm 0.139$}}}$ $2.653{\text{{$\pm 0.022$}}}$ $\mathbf{0.551}{\text{{$\pm 0.002$}}}$ $\mathbf{125.685}{\text{{$\pm 0.151$}}}$
FIM-PP (zs) (NLL) $64.281$ $\pm 0.077$ $3.949$ $\pm 0.010$ $1.988$ $\pm 0.006$ $169.687$ $\pm 0.089$
FIM-PP (zs) (S-MAPE) $65.088$ $\pm 0.037$ $3.251$ $\pm 0.007$ $4.042$ $\pm 0.025$ $174.814$ $\pm 0.102$
FIM-PP (f) (NLL) $60.106$ $\pm 0.464$ $2.428$ $\pm 0.005$ $16.068$ $\pm 0.109$ $152.528$ $\pm 0.377$
FIM-PP (f) (S-MAPE) $58.900$ $\pm 0.436$ $\mathbf{2.347}$ $\pm 0.014$ $14.734$ $\pm 0.120$ $152.547$ $\pm 0.406$
StackOverflow HYPRO $42.359{\text{{$\pm 0.170$}}}$ $1.140{\text{{$\pm 0.014$}}}$ $554{\text{{$\pm 0.010$}}} 11$$10.988{\text{{$\pm 0.559$}}}$
Dual-TPP $41.752{\text{{$\pm 0.200$}}}$ $\mathbf{1.134}{\text{{$\pm 0.019$}}}$ $514{\text{{$\pm 0.017$}}} 11$$17.582{\text{{$\pm 0.420$}}}$
A-NHP $42.591{\text{{$\pm 0.408$}}}$ $1.142{\text{{$\pm 0.011$}}}$ $1.340{\text{{$\pm 0.006$}}}$ $108.542{\text{{$\pm 0.531$}}}$
NHP $43.791{\text{{$\pm 0.147$}}}$ $1.244{\text{{$\pm 0.030$}}}$ $487{\text{{$\pm 0.004$}}} 11$$16.952{\text{{$\pm 0.404$}}}$
IFTPP $46.280{\text{{$\pm 0.892$}}}$ $1.447{\text{{$\pm 0.057$}}}$ $669{\text{{$\pm 0.005$}}} 11$$15.122{\text{{$\pm 0.627$}}}$
TCDDM $42.128{\text{{$\pm 0.591$}}}$ $1.467{\text{{$\pm 0.014$}}}$ $1.315{\text{{$\pm 0.004$}}}$ $107.659{\text{{$\pm 0.934$}}}$
CDiff $41.245{\text{{$\pm 1.400$}}}$ $141{\text{{$\pm 0.007$}}} 1.1$$199{\text{{$\pm 0.006$}}}$ $106.175{\text{{$\pm 0.340$}}}$
FIM-PP (zs) (NLL) $49.259$ $\pm 0.056$ $2.393$ $\pm 0.015$ $1.068$ $\pm 0.002$ $96.364$ $\pm 0.048$
FIM-PP (zs) (S-MAPE) $43.630$ $\pm 0.242$ $1.488$ $\pm 0.010$ $1.050$ $\pm 0.002$ $93.544$ $\pm 0.485$
FIM-PP (f) (NLL) $39.792$ $\pm 0.042$ $1.336$ $\pm 0.030$ $1.018$ $\pm 0.003$ $\mathbf{88.248}$ $\pm 0.189$
FIM-PP (f) (S-MAPE) $\mathbf{39.245}$ $\pm 0.056$ $1.407$ $\pm 0.022$ $\mathbf{1.008}$ $\pm 0.003$ $\mathbf{88.277}$ $\pm 0.231$
Amazon HYPRO $38.613{\text{{$\pm 0.536$}}}$ $\mathbf{2.007}{\text{{$\pm 0.054$}}}$ $0.477{\text{{$\pm 0.010$}}}$ $82.506{\text{{$\pm 0.840$}}}$
Dual-TPP $42.646{\text{{$\pm 0.752$}}}$ $2.562{\text{{$\pm 0.202$}}}$ $0.482{\text{{$\pm 0.012$}}}$ $86.453{\text{{$\pm 2.044$}}}$
A-NHP $39.480{\text{{$\pm 0.326$}}}$ $2.166{\text{{$\pm 0.026$}}}$ $0.476{\text{{$\pm 0.033$}}}$ $84.323{\text{{$\pm 1.815$}}}$
NHP $42.571{\text{{$\pm 0.293$}}}$ $2.561{\text{{$\pm 0.060$}}}$ $0.519{\text{{$\pm 0.023$}}}$ $92.053{\text{{$\pm 1.553$}}}$
IFTPP $43.820{\text{{$\pm 0.232$}}}$ $3.050{\text{{$\pm 0.286$}}}$ $0.481{\text{{$\pm 0.145$}}}$ $90.910{\text{{$\pm 1.611$}}}$
TCDDM $42.245{\text{{$\pm 0.174$}}}$ $2.998{\text{{$\pm 0.115$}}}$ $0.476{\text{{$\pm 0.111$}}}$ $83.826{\text{{$\pm 1.508$}}}$
CDiff $37.728{\text{{$\pm 0.199$}}}$ $2.091{\text{{$\pm 0.163$}}}$ $0.464{\text{{$\pm 0.086$}}}$ $81.987{\text{{$\pm 1.905$}}}$
FIM-PP (zs) (NLL) $46.219$ $\pm 0.108$ $2.073$ $\pm 0.012$ $0.464$ $\pm 0.001$ $128.635$ $\pm 0.398$
FIM-PP (zs) (S-MAPE) $45.153$ $\pm 0.092$ $2.106$ $\pm 0.009$ $0.453$ $\pm 0.002$ $124.476$ $\pm 0.494$
FIM-PP (f) (NLL) $\mathbf{37.208}$ $\pm 0.098$ $\mathbf{2.030}$ $\pm 0.019$ $\mathbf{0.366}$ $\pm 0.001$ $\mathbf{81.188}$ $\pm 0.142$
FIM-PP (f) (S-MAPE) $37.454$ $\pm 0.060$ $2.116$ $\pm 0.002$ $\mathbf{0.361}$ $\pm 0.002$ $87.257$ $\pm 0.198$
Retweet HYPRO $61.031{\text{{$\pm 0.092$}}}$ $2.623{\text{{$\pm 0.036$}}}$ $30.100{\text{{$\pm 0.413$}}}$ $106.110{\text{{$\pm 1.505$}}}$
Dual-TPP $61.095{\text{{$\pm 0.101$}}}$ $2.679{\text{{$\pm 0.026$}}}$ $28.914{\text{{$\pm 0.300$}}}$ $106.900{\text{{$\pm 1.293$}}}$
A-NHP $60.634{\text{{$\pm 0.097$}}}$ $2.561{\text{{$\pm 0.054$}}}$ $28.812{\text{{$\pm 0.272$}}}$ $107.234{\text{{$\pm 1.293$}}}$
NHP $60.953{\text{{$\pm 0.079$}}}$ $2.651{\text{{$\pm 0.045$}}}$ $27.130{\text{{$\pm 0.224$}}}$ $107.075{\text{{$\pm 1.398$}}}$
IFTPP $61.715{\text{{$\pm 0.152$}}}$ $2.776{\text{{$\pm 0.043$}}}$ $27.582{\text{{$\pm 0.191$}}}$ $106.711{\text{{$\pm 1.615$}}}$
TCDDM $60.501{\text{{$\pm 0.087$}}}$ $2.387{\text{{$\pm 0.050$}}}$ $27.303{\text{{$\pm 0.152$}}}$ $106.048{\text{{$\pm 0.610$}}}$
CDiff $60.661{\text{{$\pm 0.101$}}}$ $\mathbf{2.293}{\text{{$\pm 0.034$}}}$ $27.101{\text{{$\pm 0.113$}}}$ $106.184{\text{{$\pm 1.121$}}}$
FIM-PP (zs) (NLL) $60.238$ $\pm 0.161$ $4.172$ $\pm 0.064$ $24.057$ $\pm 0.050$ $99.069$ $\pm 0.390$
FIM-PP (zs) (S-MAPE) $59.392$ $\pm 0.149$ $4.323$ $\pm 0.042$ $24.804$ $\pm 0.014$ $106.317$ $\pm 0.187$
FIM-PP (f) (NLL) $59.437$ $\pm 0.082$ $2.703$ $\pm 0.012$ $21.985$ $\pm 0.014$ $\mathbf{87.585}$ $\pm 0.171$
FIM-PP (f) (S-MAPE) $\mathbf{59.150}$ $\pm 0.061$ $3.081$ $\pm 0.020$ $\mathbf{21.800}$ $\pm 0.025$ $\mathbf{87.754}$ $\pm 0.109$
Table 4: Prediction of $10$ events in test sequences of five real-world datasets. Error-bars indicate the standard deviation over $10$ trials. Results for the baseline methods were extracted from Zeng et al. (2024). Best results are bold.
[…]
Taxi HYPRO $11.875{\text{{$\pm 0.172$}}}$ $0.764{\text{{$\pm 0.008$}}}$ $0.363{\text{{$\pm 0.002$}}}$ $89.524{\text{{$\pm 0.552$}}}$
Dual-TPP $13.058{\text{{$\pm 0.220$}}}$ $0.966{\text{{$\pm 0.011$}}}$ $0.395{\text{{$\pm 0.003$}}}$ $90.812{\text{{$\pm 0.497$}}}$
A-NHP $12.542{\text{{$\pm 0.336$}}}$ $0.823{\text{{$\pm 0.007$}}}$ $0.376{\text{{$\pm 0.003$}}}$ $92.812{\text{{$\pm 0.129$}}}$
NHP $13.377{\text{{$\pm 0.184$}}}$ $0.922{\text{{$\pm 0.009$}}}$ $0.397{\text{{$\pm 0.005$}}}$ $92.182{\text{{$\pm 0.384$}}}$
IFTPP $12.765{\text{{$\pm 0.106$}}}$ $1.004{\text{{$\pm 0.013$}}}$ $0.383{\text{{$\pm 0.015$}}}$ $93.120{\text{{$\pm 0.526$}}}$
TCDDM $1.885{\text{{$\pm 0.149$}}} 1.1$$121{\text{{$\pm 0.072$}}}$ $0.385{\text{{$\pm 0.009$}}}$ $90.703{\text{{$\pm 0.356$}}}$
CDiff $11.004{\text{{$\pm 0.191$}}}$ $0.785{\text{{$\pm 0.007$}}}$ $0.350{\text{{$\pm 0.002$}}}$ $90.721{\text{{$\pm 0.291$}}}$
FIM-PP (zs) (NLL) $13.820$ $\pm 0.124$ $1.190$ $\pm 0.013$ $0.281$ $\pm 0.001$ $78.141$ $\pm 0.414$
FIM-PP (zs) (S-MAPE) $19.425$ $\pm 0.127$ $1.959$ $\pm 0.015$ $0.299$ $\pm 0.001$ $90.510$ $\pm 0.267$
FIM-PP (f) (NLL) $8.336$ $\pm 0.071$ $\mathbf{0.451}$ $\pm 0.006$ $0.291$ $\pm 0.004$ $\mathbf{75.366}$ $\pm 0.160$
FIM-PP (f) (S-MAPE) $\mathbf{8.175}$ $\pm 0.052$ $0.465$ $\pm 0.004$ $\mathbf{0.275}$ $\pm 0.001$ $\mathbf{73.529}$ $\pm 0.138$
Taobao HYPRO $21.547{\text{{$\pm 0.138$}}}$ $1.527{\text{{$\pm 0.035$}}}$ $0.591{\text{{$\pm 0.019$}}}$ $133.147{\text{{$\pm 0.341$}}}$
Dual-TPP $23.691{\text{{$\pm 0.203$}}}$ $2.674{\text{{$\pm 0.032$}}}$ $0.873{\text{{$\pm 0.010$}}}$ $139.271{\text{{$\pm 0.348$}}}$
A-NHP $21.683{\text{{$\pm 0.215$}}}$ $1.514{\text{{$\pm 0.015$}}}$ $0.608{\text{{$\pm 0.011$}}}$ $135.271{\text{{$\pm 0.395$}}}$
NHP $24.068{\text{{$\pm 0.331$}}}$ $2.769{\text{{$\pm 0.033$}}}$ $0.855{\text{{$\pm 0.013$}}}$ $137.693{\text{{$\pm 0.225$}}}$
IFTPP $23.195{\text{{$\pm 0.039$}}}$ $2.429{\text{{$\pm 0.045$}}}$ $0.602{\text{{$\pm 0.037$}}}$ $127.411{\text{{$\pm 0.573$}}}$
TCDDM $\mathbf{21.012}{\text{{$\pm 0.520$}}}$ $2.598{\text{{$\pm 0.047$}}}$ $0.610{\text{{$\pm 0.022$}}}$ $132.711{\text{{$\pm 0.774$}}}$
CDiff $\mathbf{21.221}{\text{{$\pm 0.176$}}}$ $1.416{\text{{$\pm 0.024$}}}$ $\mathbf{0.535}{\text{{$\pm 0.016$}}}$ $\mathbf{126.824}{\text{{$\pm 0.366$}}}$
FIM-PP (zs) (NLL) $31.880$ $\pm 0.040$ $2.024$ $\pm 0.004$ $1.955$ $\pm 0.011$ $170.278$ $\pm 0.029$
FIM-PP (zs) (S-MAPE) $32.249$ $\pm 0.041$ $1.822$ $\pm 0.004$ $3.538$ $\pm 0.039$ $172.683$ $\pm 0.194$
FIM-PP (f) (NLL) $27.974$ $\pm 0.162$ $\mathbf{1.325}$ $\pm 0.010$ $14.954$ $\pm 0.253$ $145.821$ $\pm 1.120$
FIM-PP (f) (S-MAPE) $27.940$ $\pm 0.075$ $1.358$ $\pm 0.009$ $14.844$ $\pm 0.055$ $153.499$ $\pm 0.553$
StackOverflow HYPRO $21.062{\text{{$\pm 0.372$}}}$ $0.921{\text{{$\pm 0.019$}}}$ $1.235{\text{{$\pm 0.006$}}}$ $107.566{\text{{$\pm 0.218$}}}$
Dual-TPP $21.229{\text{{$\pm 0.394$}}}$ $0.936{\text{{$\pm 0.013$}}}$ $1.223{\text{{$\pm 0.010$}}}$ $107.274{\text{{$\pm 0.200$}}}$
A-NHP $22.019{\text{{$\pm 0.220$}}}$ $0.978{\text{{$\pm 0.023$}}}$ $1.225{\text{{$\pm 0.007$}}}$ $100.137{\text{{$\pm 0.167$}}}$
NHP $21.655{\text{{$\pm 0.314$}}}$ $0.970{\text{{$\pm 0.014$}}}$ $1.266{\text{{$\pm 0.003$}}}$ $108.867{\text{{$\pm 0.361$}}}$
IFTPP $22.339{\text{{$\pm 0.322$}}}$ $0.970{\text{{$\pm 0.011$}}}$ $1.251{\text{{$\pm 0.005$}}}$ $105.674{\text{{$\pm 0.337$}}}$
TCDDM $22.042{\text{{$\pm 0.193$}}}$ $1.205{\text{{$\pm 0.014$}}}$ $1.228{\text{{$\pm 0.010$}}}$ $108.111{\text{{$\pm 0.112$}}}$
CDiff $20.191{\text{{$\pm 0.455$}}}$ $0.916{\text{{$\pm 0.010$}}}$ $1.180{\text{{$\pm 0.003$}}}$ $102.367{\text{{$\pm 0.267$}}}$
FIM-PP (zs) (NLL) $23.527$ $\pm 0.033$ $1.188$ $\pm 0.005$ $1.039$ $\pm 0.003$ $92.919$ $\pm 0.556$
FIM-PP (zs) (S-MAPE) $21.182$ $\pm 0.146$ $0.857$ $\pm 0.010$ $1.047$ $\pm 0.002$ $93.715$ $\pm 0.388$
FIM-PP (f) (NLL) $\mathbf{19.938}$ $\pm 0.093$ $\mathbf{0.823}$ $\pm 0.010$ $\mathbf{1.012}$ $\pm 0.004$ $\mathbf{87.503}$ $\pm 0.402$
FIM-PP (f) (S-MAPE) $\mathbf{19.846}$ $\pm 0.152$ $\mathbf{0.832}$ $\pm 0.010$ $\mathbf{1.004}$ $\pm 0.003$ $\mathbf{87.110}$ $\pm 0.393$
Amazon HYPRO $24.956{\text{{$\pm 0.663$}}}$ $1.765{\text{{$\pm 0.039$}}}$ $0.442{\text{{$\pm 0.015$}}}$ $83.401{\text{{$\pm 1.033$}}}$
Dual-TPP $25.929{\text{{$\pm 0.280$}}}$ $2.098{\text{{$\pm 0.101$}}}$ $0.475{\text{{$\pm 0.008$}}}$ $82.352{\text{{$\pm 1.285$}}}$
A-NHP $24.116{\text{{$\pm 0.807$}}}$ $1.741{\text{{$\pm 0.039$}}}$ $0.454{\text{{$\pm 0.014$}}}$ $84.323{\text{{$\pm 1.815$}}}$
NHP $25.730{\text{{$\pm 0.497$}}}$ $1.843{\text{{$\pm 0.053$}}}$ $0.491{\text{{$\pm 0.048$}}}$ $89.135{\text{{$\pm 1.092$}}}$
IFTPP $26.632{\text{{$\pm 0.519$}}}$ $1.955{\text{{$\pm 0.112$}}}$ $0.464{\text{{$\pm 0.066$}}}$ $89.305{\text{{$\pm 1.288$}}}$
TCDDM $25.091{\text{{$\pm 0.227$}}}$ $1.778{\text{{$\pm 0.090$}}}$ $0.448{\text{{$\pm 0.082$}}}$ $\mathbf{82.105}{\text{{$\pm 1.564$}}}$
CDiff $24.230{\text{{$\pm 0.287$}}}$ $1.766{\text{{$\pm 0.079$}}}$ $0.450{\text{{$\pm 0.049$}}}$ $\mathbf{82.124}{\text{{$\pm 2.094$}}}$
FIM-PP (zs) (NLL) $21.736$ $\pm 0.115$ $1.141$ $\pm 0.010$ $0.449$ $\pm 0.002$ $120.894$ $\pm 0.393$
FIM-PP (zs) (S-MAPE) $21.418$ $\pm 0.142$ $1.198$ $\pm 0.008$ $0.445$ $\pm 0.002$ $118.944$ $\pm 0.487$
FIM-PP (f) (NLL) $\mathbf{18.428}$ $\pm 0.124$ $\mathbf{1.091}$ $\pm 0.016$ $0.361$ $\pm 0.001$ $87.264$ $\pm 0.323$
FIM-PP (f) (S-MAPE) $\mathbf{18.566}$ $\pm 0.072$ $1.153$ $\pm 0.007$ $\mathbf{0.352}$ $\pm 0.002$ $\mathbf{82.555}$ $\pm 0.278$
Retweet HYPRO $31.743{\text{{$\pm 0.068$}}}$ $1.927{\text{{$\pm 0.027$}}}$ $33.683{\text{{$\pm 0.245$}}}$ $105.073{\text{{$\pm 0.958$}}}$
Dual-TPP $31.652{\text{{$\pm 0.075$}}}$ $1.963{\text{{$\pm 0.038$}}}$ $28.104{\text{{$\pm 0.486$}}}$ $106.721{\text{{$\pm 0.774$}}}$
A-NHP $\mathbf{30.337}{\text{{$\pm 0.065$}}}$ $1.823{\text{{$\pm 0.031$}}}$ $26.310{\text{{$\pm 0.333$}}}$ $106.021{\text{{$\pm 1.011$}}}$
NHP $30.817{\text{{$\pm 0.090$}}}$ $1.713{\text{{$\pm 0.024$}}}$ $27.010{\text{{$\pm 0.429$}}}$ $107.053{\text{{$\pm 1.390$}}}$
IFTPP $31.974{\text{{$\pm 0.032$}}}$ $1.942{\text{{$\pm 0.062$}}}$ $28.825{\text{{$\pm 0.221$}}}$ $106.014{\text{{$\pm 0.633$}}}$
TCDDM $32.006{\text{{$\pm 0.074$}}}$ $1.789{\text{{$\pm 0.094$}}}$ $29.124{\text{{$\pm 0.405$}}}$ $106.738{\text{{$\pm 0.791$}}}$
CDiff $31.237{\text{{$\pm 0.078$}}}$ $1.745{\text{{$\pm 0.036$}}}$ $26.429{\text{{$\pm 0.201$}}}$ $105.767{\text{{$\pm 0.771$}}}$
FIM-PP (zs) (NLL) $31.027$ $\pm 0.031$ $2.355$ $\pm 0.032$ $27.085$ $\pm 0.002$ $97.590$ $\pm 0.152$
FIM-PP (zs) (S-MAPE) $30.986$ $\pm 0.104$ $2.396$ $\pm 0.012$ $28.305$ $\pm 0.056$ $105.397$ $\pm 0.496$
FIM-PP (f) (NLL) $30.592$ $\pm 0.037$ $\mathbf{1.611}$ $\pm 0.031$ $25.021$ $\pm 0.034$ $86.875$ $\pm 0.108$
FIM-PP (f) (S-MAPE) $30.788$ $\pm 0.051$ $1.652$ $\pm 0.018$ $\mathbf{24.836}$ $\pm 0.042$ $\mathbf{85.222}$ $\pm 0.086$
Table 5: Prediction of $5$ events in test sequences of five real-world datasets. Error-bars indicate the standard deviation over $10$ trials. Results for the baseline methods were extracted from Zeng et al. (2024). Best results are bold.
[…]
Taxi HYPRO $5.952{\text{{$\pm 0.126$}}}$ $0.500{\text{{$\pm 0.011$}}}$ $0.322{\text{{$\pm 0.004$}}}$ $85.994{\text{{$\pm 0.227$}}}$
Dual-TPP $7.534{\text{{$\pm 0.111$}}}$ $0.636{\text{{$\pm 0.009$}}}$ $0.340{\text{{$\pm 0.003$}}}$ $89.727{\text{{$\pm 0.320$}}}$
A-NHP $6.441{\text{{$\pm 0.090$}}}$ $0.682{\text{{$\pm 0.010$}}}$ $0.347{\text{{$\pm 0.002$}}}$ $89.070{\text{{$\pm 0.152$}}}$
NHP $7.405{\text{{$\pm 0.122$}}}$ $0.641{\text{{$\pm 0.013$}}}$ $0.351{\text{{$\pm 0.008$}}}$ $91.625{\text{{$\pm 0.177$}}}$
IFTPP $7.209{\text{{$\pm 0.184$}}}$ $0.608{\text{{$\pm 0.008$}}}$ $0.335{\text{{$\pm 0.003$}}}$ $90.512{\text{{$\pm 0.169$}}}$
TCDDM $5.877{\text{{$\pm 0.095$}}}$ $0.648{\text{{$\pm 0.015$}}}$ $0.327{\text{{$\pm 0.005$}}}$ $88.051{\text{{$\pm 0.240$}}}$
CDiff $5.966{\text{{$\pm 0.083$}}}$ $0.547{\text{{$\pm 0.007$}}}$ $0.318{\text{{$\pm 0.003$}}}$ $89.535{\text{{$\pm 0.294$}}}$
FIM-PP (zs) (NLL) $6.773$ $\pm 0.064$ $0.655$ $\pm 0.013$ $\mathbf{0.246}$ $\pm 0.001$ $74.912$ $\pm 0.793$
FIM-PP (zs) (SMAPE) $6.763$ $\pm 0.018$ $0.654$ $\pm 0.005$ $0.248$ $\pm 0.001$ $76.402$ $\pm 0.802$
FIM-PP (f) (NLL) $\mathbf{4.083}$ $\pm 0.032$ $\mathbf{0.311}$ $\pm 0.007$ $0.250$ $\pm 0.002$ $\mathbf{71.108}$ $\pm 0.902$
FIM-PP (f) (SMAPE) $4.103$ $\pm 0.029$ $\mathbf{0.315}$ $\pm 0.005$ $0.248$ $\pm 0.001$ $73.181$ $\pm 0.409$
Taobao HYPRO $11.317{\text{{$\pm 0.111$}}}$ $0.817{\text{{$\pm 0.037$}}}$ $0.573{\text{{$\pm 0.011$}}}$ $133.837{\text{{$\pm 0.524$}}}$
Dual-TPP $13.280{\text{{$\pm 0.092$}}}$ $1.877{\text{{$\pm 0.014$}}}$ $0.691{\text{{$\pm 0.007$}}}$ $134.437{\text{{$\pm 0.458$}}}$
A-NHP $11.223{\text{{$\pm 0.145$}}}$ $0.873{\text{{$\pm 0.023$}}}$ $0.550{\text{{$\pm 0.014$}}}$ $132.266{\text{{$\pm 0.532$}}}$
NHP $11.973{\text{{$\pm 0.176$}}}$ $1.910{\text{{$\pm 0.031$}}}$ $0.712{\text{{$\pm 0.017$}}}$ $134.693{\text{{$\pm 0.369$}}}$
IFTPP $11.052{\text{{$\pm 0.108$}}}$ $1.941{\text{{$\pm 0.049$}}}$ $0.601{\text{{$\pm 0.017$}}}$ $126.320{\text{{$\pm 0.591$}}}$
TCDDM $11.609{\text{{$\pm 0.184$}}}$ $1.690{\text{{$\pm 0.023$}}}$ $0.675{\text{{$\pm 0.009$}}}$ $129.009{\text{{$\pm 0.923$}}}$
CDiff $\mathbf{10.147}{\text{{$\pm 0.140$}}}$ $\mathbf{0.730}{\text{{$\pm 0.019$}}}$ $\mathbf{0.519}{\text{{$\pm 0.008$}}}$ $\mathbf{124.339}{\text{{$\pm 0.322$}}}$
FIM-PP (zs) (NLL) $15.951$ $\pm 0.042$ $1.129$ $\pm 0.007$ $1.761$ $\pm 0.013$ $168.299$ $\pm 0.249$
FIM-PP (zs) (SMAPE) $15.955$ $\pm 0.034$ $1.106$ $\pm 0.007$ $1.918$ $\pm 0.026$ $167.486$ $\pm 0.199$
FIM-PP (f) (NLL) $13.173$ $\pm 0.261$ $0.745$ $\pm 0.010$ $14.892$ $\pm 0.370$ $146.921$ $\pm 0.858$
FIM-PP (f) (SMAPE) $13.572$ $\pm 0.166$ $0.759$ $\pm 0.005$ $17.384$ $\pm 0.416$ $150.562$ $\pm 0.830$
StackOverflow HYPRO $11.590{\text{{$\pm 0.186$}}}$ $0.586{\text{{$\pm 0.019$}}}$ $1.227{\text{{$\pm 0.018$}}}$ $109.014{\text{{$\pm 0.422$}}}$
Dual-TPP $11.719{\text{{$\pm 0.109$}}}$ $0.591{\text{{$\pm 0.026$}}}$ $1.296{\text{{$\pm 0.010$}}}$ $106.697{\text{{$\pm 0.381$}}}$
A-NHP $11.595{\text{{$\pm 0.197$}}}$ $0.575{\text{{$\pm 0.009$}}}$ $1.188{\text{{$\pm 0.014$}}}$ $105.799{\text{{$\pm 0.516$}}}$
NHP $11.807{\text{{$\pm 0.155$}}}$ $0.596{\text{{$\pm 0.015$}}}$ $1.261{\text{{$\pm 0.013$}}}$ $108.074{\text{{$\pm 0.661$}}}$
IFTPP $13.124{\text{{$\pm 0.174$}}}$ $0.702{\text{{$\pm 0.008$}}}$ $1.182{\text{{$\pm 0.039$}}}$ $108.409{\text{{$\pm 0.692$}}}$
TCDDM $11.410{\text{{$\pm 0.129$}}}$ $0.630{\text{{$\pm 0.015$}}}$ $1.201{\text{{$\pm 0.028$}}}$ $107.893{\text{{$\pm 0.942$}}}$
CDiff $10.735{\text{{$\pm 0.183$}}}$ $0.571{\text{{$\pm 0.012$}}}$ $1.153{\text{{$\pm 0.011$}}}$ $100.586{\text{{$\pm 0.299$}}}$
FIM-PP (zs) (NLL) $11.520$ $\pm 0.057$ $0.657$ $\pm 0.003$ $1.030$ $\pm 0.001$ $93.296$ $\pm 0.506$
FIM-PP (zs) (SMAPE) $11.334$ $\pm 0.055$ $0.631$ $\pm 0.004$ $1.020$ $\pm 0.000$ $91.864$ $\pm 0.443$
FIM-PP (f) (NLL) $\mathbf{10.353}$ $\pm 0.051$ $\mathbf{0.527}$ $\pm 0.004$ $0.990$ $\pm 0.003$ $86.443$ $\pm 0.128$
FIM-PP (f) (SMAPE) $\mathbf{10.341}$ $\pm 0.013$ $\mathbf{0.525}$ $\pm 0.007$ $\mathbf{0.984}$ $\pm 0.003$ $\mathbf{86.133}$ $\pm 0.212$
Amazon HYPRO $9.552{\text{{$\pm 0.172$}}}$ $1.397{\text{{$\pm 0.033$}}}$ $0.433{\text{{$\pm 0.008$}}}$ $82.847{\text{{$\pm 0.748$}}}$
Dual-TPP $11.309{\text{{$\pm 0.093$}}}$ $1.742{\text{{$\pm 0.302$}}}$ $0.476{\text{{$\pm 0.010$}}}$ $86.633{\text{{$\pm 0.573$}}}$
A-NHP $\mathbf{9.430}{\text{{$\pm 0.131$}}}$ $1.117{\text{{$\pm 0.049$}}}$ $0.427{\text{{$\pm 0.033$}}}$ $83.121{\text{{$\pm 0.415$}}}$
NHP $11.273{\text{{$\pm 0.198$}}}$ $1.431{\text{{$\pm 0.024$}}}$ $0.501{\text{{$\pm 0.009$}}}$ $90.591{\text{{$\pm 0.667$}}}$
IFTPP $10.230{\text{{$\pm 0.224$}}}$ $1.663{\text{{$\pm 0.168$}}}$ $0.447{\text{{$\pm 0.015$}}}$ $88.900{\text{{$\pm 0.610$}}}$
TCDDM $10.557{\text{{$\pm 0.331$}}}$ $1.409{\text{{$\pm 0.203$}}}$ $0.460{\text{{$\pm 0.032$}}}$ $82.401{\text{{$\pm 0.810$}}}$
CDiff $\mathbf{9.478}{\text{{$\pm 0.081$}}}$ $1.326{\text{{$\pm 0.082$}}}$ $0.424{\text{{$\pm 0.018$}}}$ $81.287{\text{{$\pm 0.994$}}}$
FIM-PP (zs) (NLL) $11.124$ $\pm 0.059$ $\mathbf{0.736}$ $\pm 0.004$ $0.449$ $\pm 0.004$ $119.129$ $\pm 0.746$
FIM-PP (zs) (SMAPE) $11.029$ $\pm 0.033$ $\mathbf{0.735}$ $\pm 0.005$ $0.445$ $\pm 0.003$ $117.793$ $\pm 0.635$
FIM-PP (f) (NLL) $10.034$ $\pm 0.060$ $\mathbf{0.737}$ $\pm 0.006$ $\mathbf{0.341}$ $\pm 0.004$ $\mathbf{78.738}$ $\pm 0.339$
FIM-PP (f) (SMAPE) $10.004$ $\pm 0.019$ $\mathbf{0.732}$ $\pm 0.003$ $\mathbf{0.343}$ $\pm 0.004$ $79.623$ $\pm 0.524$
Retweet HYPRO $6.145{\text{{$\pm 0.096$}}} 1.1$$105{\text{{$\pm 0.026$}}}$ $27.236{\text{{$\pm 0.259$}}}$ $103.052{\text{{$\pm 1.206$}}}$
Dual-TPP $16.050{\text{{$\pm 0.085$}}}$ $1.077{\text{{$\pm 0.027$}}}$ $31.493{\text{{$\pm 0.162$}}}$ $101.322{\text{{$\pm 1.127$}}}$
A-NHP $16.124{\text{{$\pm 0.089$}}}$ $1.058{\text{{$\pm 0.029$}}}$ $29.247{\text{{$\pm 0.145$}}}$ $105.930{\text{{$\pm 1.380$}}}$
NHP $5.945{\text{{$\pm 0.094$}}} 1.1$$113{\text{{$\pm 0.040$}}}$ $32.367{\text{{$\pm 0.104$}}}$ $107.022{\text{{$\pm 1.077$}}}$
IFTPP $16.043{\text{{$\pm 0.222$}}}$ $1.313{\text{{$\pm 0.011$}}}$ $30.853{\text{{$\pm 0.119$}}}$ $106.941{\text{{$\pm 2.031$}}}$
TCDDM $5.874{\text{{$\pm 0.053$}}} 1.1$$194{\text{{$\pm 0.021$}}}$ $28.530{\text{{$\pm 0.110$}}}$ $105.570{\text{{$\pm 0.940$}}}$
CDiff $15.858{\text{{$\pm 0.080$}}}$ $\mathbf{1.023}{\text{{$\pm 0.036$}}}$ $26.078{\text{{$\pm 0.175$}}}$ $106.620{\text{{$\pm 1.008$}}}$
FIM-PP (zs) (NLL) $15.747$ $\pm 0.032$ $1.342$ $\pm 0.027$ $28.138$ $\pm 0.068$ $98.668$ $\pm 0.794$
FIM-PP (zs) (SMAPE) $15.780$ $\pm 0.017$ $1.331$ $\pm 0.030$ $28.472$ $\pm 0.071$ $100.765$ $\pm 0.803$
FIM-PP (f) (NLL) $15.645$ $\pm 0.020$ $\mathbf{1.033}$ $\pm 0.034$ $\mathbf{25.308}$ $\pm 0.135$ $\mathbf{83.010}$ $\pm 0.278$
FIM-PP (f) (SMAPE) $\mathbf{15.585}$ $\pm 0.040$ $\mathbf{1.012}$ $\pm 0.040$ $\mathbf{25.413}$ $\pm 0.086$ $84.945$ $\pm 0.343$
A.3 Fine-Tuning
Due to the relatively small parameter count of 16.116.1M, finetuning takes just a few minutes and less than 11GB of GPU memory for all datasets. In Figure 5 we compare the finetuning speed against training a FIM-PP from scratch, i.e. with randomly initialized parameters. The results indicate that pre-training yields a good initialization for finetuning, which converges much faster than training from randomly initialized set of parameters. Moreover, the finetuned model generally archives lower errors on the test split.
Figure 5: Comparison of the fine-tuning loss curves of a pre-trained FIM-$P$ model versus random initialization. Note that one epoch corresponds to just one inference-path prediction and is therefore very fast. Our results indicate that the pre-training achieves faster convergence as well as a higher loglikelihood when converged.
A.4 Performance with varying Context Size
We also investigated the sensitivity of FIM-PP to the number of context paths passed to the model. We studied this behavior on data from three synthetic processes and evaluated the results based on the sMAPE metric. Figure 6 contains the results of this experiment. Initially, providing more context paths to the model improves the accuracy of the estimated process. After a few hundred context paths, the performance saturates; more paths only improve marginally improve the estimate. Crucially, the performance does not degrade with more paths, highlighting the robustness of our model.
Figure 6: sMAPE error against the ground truth intensity for varying number of context paths. We used 100 points per path. Our results indicate that the performance of the model does not noticeable improve for more than 500 paths, at least for the synthetic datasets tested.
Appendix B Data Generation
To train our Foundation Inference Model, we generate a comprehensive synthetic dataset of marked temporal point processes. Each process is an instance of a multivariate Hawkes process. The conditional intensity function $\lambda(t,\kappa\mid\mathcal{H}_{t})$ at time $t\in\mathbb{R}_{+}$ for a mark $\kappa\in\mathcal{K}$ given a history of marked events $\mathcal{H}_{t}=\{(t_{i},\kappa_{i})\mid t_{i}<t\}$ is defined as
$\lambda(t,\kappa\mid\mathcal{H}_{t})=\max\left(0,\mu_{\kappa}(t)+\sum_{(t^{\prime},\kappa^{\prime})\in\mathcal{H}_{t}}z_{\kappa\kappa^{\prime}}\gamma_{\kappa\kappa^{\prime}}(t-t^{\prime})\right)\;,$ (13)
where $\mu_{\kappa}$ is the time-dependent base intensity for mark $\kappa$, $\gamma_{\kappa\kappa^{\prime}}$ is the interaction kernel between $\kappa$ and $\kappa^{\prime}\in\mathcal{K}$ and $z_{\kappa\kappa^{\prime}}\in\mathbb{R}$ are sampled pre-factors of the interaction, varying the interaction behavior further.
To the best of our knowledge, no open-source solution for sampling such Hawkes processes with time-dependent base intensity functions exists. Hence, we implemented an efficient custom sampling library for such processes in C++. We will release the source code of this library in the supplementary material of our work.
B.1 Dataset Configurations
We sample Hawkes processe instances over a set of marks $\mathcal{K}$ in equation 13 in two stages.
At first, the functional forms for the base intensities $\mu_{\kappa}$ and interaction kernels $\gamma_{\kappa\kappa^{\prime}}$ are drawn from a library of parametric functions. The parameters for these functions are then sampled from specified prior distributions. Our used functional forms and their parameters are summarized in Table 6. The parameter ranges were chosen more or less arbitrarily so that the paths look realistic. We keep these choices fixed and did not modify them based on the performance on our evaluation sets in order to prevent overfitting to those, which would be contrary to the concept of a foundation model.
The pre-factors further diversify the sampled processes by introducing sparse connectivity and inhibitory effects. For each process with interactions, we choose one of two pre-factor distributions $Z_{\text{{strong}}}$ and $Z_{\text{{sparse}}}$ on $\{-1,0,1\}$, which differ by their induced connectivity:
$Z_{\text{{strong}}}$ $=\text{Categorical}\left(-1:0.06,\;0:0.4,\;1:0.54\right)$ (14)
$Z_{\text{{sparse}}}$ $=\text{Categorical}\left(-1:0.01,\;0:0.9,\;1:0.09\right)$ (15)
In other words, for $Z_{\text{{strong}}}$, only $40\%$ of interactions will be non-influencing, while for $Z_{\text{{sparse}}}$, $90\%$ of interactions will be non-influencing. For influencing interactions, $90\%$ will be excitatory, while $10\%$ will be inhibitory.
Once the full intensity function for a process is defined, event sequences are generated using Ogata’s modified thinning algorithm.
Figure 7 contains summarizing statistics of our train distribution. Aggregated, this prior is very broad. Importantly, it covers distributions of the real-world datasets in our experiments. The corresponding statistics for these datasets are depicted in Figures 8 to 12.
Figure 7: Distribution of our synthetic training data for Hawkes processes with 22 marks. The distribution is almost uniform and very broad and therefore also captures real world phenomena.
Figure 8: Amazon dataset statistics.
Figure 9: Taxi dataset statistics.
Figure 10: Taobao dataset statistics.
Figure 11: Stackoverflow dataset statistics.
Figure 12: Retweet dataset statistics.
B.2 Dataset Size
We sample instances of every Hawkes process configuration in Table 6, and simulate them for different number of marks, sequences and events, detailed in Table 7. In total, our training data consisted of 7272k processes and 14.414.4M events.
Table 6: Summary of the parametrized base intensities and interaction kernels of Hawkes processes used in our synthetic data generation. The parameters of each configuration are sampled from uniform distributions, covering a wide range of processes.
Dataset Configuration Base Intensity $\mu(t)$ Interaction Kernel $\gamma(t)$ Parameter Distributions
Constant Base & Constant Exponential Decay $c_{0}\sim\mathcal{U}(0.01,1.3)$
Exponential Kernel $\mu(t)=c_{0}$ $\gamma(t)=\alpha e^{-\beta t}$ $\alpha\sim\mathcal{U}(0.005,1.0)$
(no interactions) $(\gamma_{ij,\,i\neq j}=0)$ $\beta\sim\mathcal{U}(0.001,10.0)$
Constant Base & Constant Exponential Decay $c_{0}\sim\mathcal{U}(0.01,1.3)$
Exponential Kernel $\mu(t)=c_{0}$ $\gamma(t)=\alpha e^{-\beta t}$ $\alpha\sim\mathcal{U}(0.005,1.0)$
$\beta\sim\mathcal{U}(0.001,10.0)$
Sinusoidal Base & Sinusoidal Exponential Decay $c_{0}\sim\mathcal{U}(0.05,0.15)$
Exponential Kernel $\mu(t)=A\sin(\omega(t-\gamma))+c_{0}$ $\gamma(t)=\alpha e^{-\beta t}$ $A\sim\mathcal{U}(0.0,10.0)$
$\omega\sim\mathcal{U}(0.1,15.0)$
$\gamma\sim\mathcal{U}(0.0,5.0)$
$\alpha\sim\mathcal{U}(0.1,0.6)$
$\beta\sim\mathcal{U}(0.8,2.0)$
Gamma Base & Gamma Shape + Constant Exponential Decay $c_{0}\sim\mathcal{U}(0.1,1.3)$
Exponential Kernel $\mu(t)=At^{p}e^{-\beta_{0}t}+c_{0}$ $\gamma(t)=\alpha e^{-\beta_{1}t}$ $A\sim\mathcal{U}(10.0,50.0)$
$p\sim\mathcal{U}(1.0,2.0)$
$\beta_{0}\sim\mathcal{U}(1.0,10.1)$
$\alpha\sim\mathcal{U}(0.005,1.0)$
$\beta_{1}\sim\mathcal{U}(0.001,10.0)$
Poisson Process Constant Zero Kernel $c_{0}\sim\mathcal{U}(0.01,1.3)$
$\mu(t)=c_{0}$ $\gamma(t)=0$
Constant Base & Constant Rayleigh $c_{0}\sim\mathcal{U}(0.01,1.3)$
Rayleigh Kernel $\mu(t)=c_{0}$ $\gamma(t)=a_{0}\frac{(t-t_{\text{shift}})}{a_{1}^{2}}\exp\!\big(-\tfrac{(t-t_{\text{shift}})^{2}}{2a_{1}^{2}}\big)\,$ $a_{0}\sim\mathcal{U}(0.001,1.0)$
$a_{1}\sim\mathcal{U}(0.05,0.25)$
$t_{\text{shift}}\sim\mathcal{U}(0.0,0.1)$
Table 7: For each dataset configuration, we sample Hawkes processes with varying numbers (#\#) of marks, sequences and events per sequence.
#\# Marks #\# Samples #\# Sequences #\# Events per Sequence
11 $1000$ $2000$ $100$
55 $1000$ $2000$ $100$
$10$ $1000$ $2000$ $100$
$15$ $1000$ $2000$ $100$
$22$ $5000$ $2000$ $100$
Appendix C Instance Normalization
To ensure that FIM-PP can generalize across datasets with vastly different time scales, we introduce an instance normalization scheme that makes the model agnostic to the absolute units of time. Let $\mathcal{C}=\{\mathcal{S}^{j}\}_{j=1}^{m}$ denote a context of FIM-PP, i.e. a set of marked event sequences $\mathcal{S}^{j}=\{(t^{j}_{i},\kappa^{j}_{i})\}_{i=1}^{n_{j}}$. Identifying $t_{0}^{j}=0$, we define the inter-event times as $\Delta t_{i}^{j}=t_{i}^{j}-t_{i-1}^{j}$ and the maximum inter-event time in the context as
$\Delta t_{\text{{max}}}^{\text{{cont}}}=\max_{j=1,\dots,m}\max_{i=1,\dots,n_{j}}\Delta t_{i}^{j}\quad.$ (16)
All time-related inputs to the model, including context event times $t_{i}^{j}$, inter-event times features $\Delta t_{i}^{j}$, and history event times (e.g. from a target sequence during training) $t$, are scaled by the maximum inter-event time:
$t^{\prime}=\frac{t}{\Delta t_{\text{{max}}}^{\text{{cont}}}}.$ (17)
This transformation maps all temporal information to a canonical scale where the largest inter-event gap becomes $1$.
This change of time variable also transforms the intensity function. To preserve the number of expected events within a differential interval, the intensities must be related by $\lambda(t)dt=\lambda^{\prime}(t^{\prime})dt^{\prime}$. Since $dt=\Delta t_{\text{{max}}}^{\text{{cont}}}\,dt^{\prime}$, it follows that the intensity in the normalized time domain, $\lambda^{\prime}(t^{\prime})$, is a scaled version of the original:
$\lambda^{\prime}(t^{\prime})=\Delta t_{\text{{max}}}^{\text{{cont}}}\cdot\lambda(t).$ (18)
Consequently, the model is trained to predict this normalized intensity $\lambda^{\prime}(t^{\prime})$. During inference, to obtain the intensity in the original, real-world time scale, the model’s output is simply denormalized by dividing by the same constant $\Delta t_{\text{{max}}}^{\text{{cont}}}$. This entire process allows the FIM to learn scale-invariant temporal dynamics, a key requirement for effective zero-shot inference on unseen data.
Appendix D Training Details
Our Foundation Inference Model was trained on the comprehensive synthetic datasets described in Appendix B. The training took about 5 days on a single NVIDIA A100-80GB.
D.1 Data Handling and Batching
Each sample in our dataset represents a single underlying process, comprising up to 2000 distinct time series paths. During training, we dynamically partition these paths into context and inference sets for each batch.
On-the-fly Path Selection
For each sample, we randomly select a single path ($P_{\text{inference}}=1$) to serve as the inference target. The remaining paths are designated as the context set. To train a model that is robust to varying amounts of contextual information, the number of context paths presented in each training step is randomized. Specifically, for each sample in a batch, we uniformly sample a number of context paths between a minimum of 400 and a maximum of 2000.
Variable Sequence Lengths
As a form of data augmentation, we also vary the length of the historical sequences. For 90% of the training batches, all sequences (both context and inference) are truncated to a random length chosen uniformly from the interval $[15,100]$. For the remaining 10% of batches, the full sequence length of 100 events is used. This strategy encourages the model to make reliable predictions from both short and long historical contexts. For validation, we use fixed, full-length sequences to ensure consistent and comparable evaluation metrics.
D.2 Hyperparameters and Optimization
The model architecture is based on the Transformer Vaswani et al. (2017). Context sequences are processed by a 4-layer Transformer encoder, and the resulting path summaries are further refined by a 2-layer Transformer encoder. The history of the target sequence is processed by a 4-layer Transformer decoder, which attends to the context summary as memory. Both encoders and the decoder use 4 attention heads and a hidden dimension of 256. The final intensity parameters ($\mu,\alpha,\beta$) are predicted by three separate Multi-Layer Perceptrons (MLPs), each with two hidden layers of 256 units.
In total, our model has $16.1$ million trainable parameters.
We trained the model using the AdamW optimizer Loshchilov and Hutter (2019) with a learning rate of $5\times 10^{-5}$ and a weight decay of $10^{-4}$. To accelerate computation, we utilized bfloat16 mixed-precision training.
D.3 Train Objective
We use the standard negative log-likelihood (NLL) for a marked temporal point process as the train objective for FIM-PP. By Section 3, the MTPP density at a sequence of events $\mathcal{S}=\{(t_{i},\kappa_{i})\}_{i=1}^{n}$ in the interval $[0,T]$ is
$f\left(\{(t_{i},\kappa_{i}\}_{i=1}^{n}\right)=\left[\prod_{i=1}^{n}\lambda(t_{i},\kappa_{i}\mid\mathcal{H}_{t_{i}})\right]\exp\left(-\int_{0}^{T}\lambda(s\mid\mathcal{H}_{s})ds\right)\;.$ (19)
Thus, the NLL of a target sequence under the distribution induced the model’s predicted intensity function $\hat{\lambda}$ is
$\mathcal{L}_{\text{NLL}}=\sum_{\kappa\in\mathcal{K}}\hat{\Lambda}(T,\kappa)-\sum_{(t,\kappa)\in\mathcal{T}}\hat{\lambda}(t,\kappa\mid\mathcal{H}_{t})\;.$ (20)
where $\hat{\Lambda}(T,\kappa)=\int_{0}^{T}\hat{\lambda}(s,\kappa\mid\mathcal{H}_{s})ds$ is the predicted integrated intensity. We approximate the integral using Monte Carlo integration
$\hat{\Lambda}(T,\kappa)\approx\frac{T}{N_{\text{MC}}}\sum_{i=1}^{N_{\text{MC}}}\hat{\lambda}(s_{i},\kappa\mid\mathcal{H}_{s_{i}}),$ (21)
with $N_{\text{MC}}=100$ samples and $s_{i}\sim\mathcal{U}(0,T)$.
Appendix E Evaluation Datasets
To evaluate the inference capabilities of FIM-PP, we use five widely-recognized real-world datasets that were not seen during training:
Amazon
This dataset comprises sequences of product reviews from users on the Amazon platform, collected over a ten-year period from 2008 to 2018 Ni et al. (2019). Each sequence represents the review history of a single user. An event is defined by the timestamp of a review, and its mark corresponds to one of 16 distinct product categories. The analysis is performed on a subset of 5,200 of the most active users to ensure sequences are sufficiently long for meaningful analysis.
Taxi
Derived from New York City’s public taxi trip records, this dataset captures the operational patterns of taxi drivers. Each sequence corresponds to the activity log of an individual driver. Events are either pick-ups or drop-offs, and the event marks are defined by the combination of the event type (pick-up/drop-off) and the borough where it occurred, resulting in 10 unique marks. The dataset consists of sequences from a random sample of 2,000 drivers.
Taobao
This dataset originates from the 2018 Tianchi Big Data Competition and contains logs of user interactions on the Taobao e-commerce platform over a period in late 2017 Zhu et al. (2018). The sequences track the behavior of anonymous users, including actions like browsing and purchasing. The 17 event types correspond to different product category groups. For the evaluation, sequences from the 2,000 most active users are utilized.
StackOverflow
Sourced from the popular question-and-answering website StackOverflow, this dataset tracks the awarding of achievement badges to users over a two-year span Leskovec and Krevl (2014). Each sequence represents a user’s history of earned badges. The events are the timestamps when badges were awarded, and the marks are the 22 different types of badges available on the platform. The evaluation subset includes 2,200 active users.
Retweet
This dataset tracks the dynamics of information spread through time-stamped user retweet sequences Zhou et al. (2013). Each sequence corresponds to the retweet history of an individual user. An event is defined by the timestamp of a retweet, and its mark is categorized into one of three types based on the influence of the original poster: ”small” (fewer than 120 followers), ”medium” (fewer than 1,363 followers), and ”large” (all other users). The analysis is performed on a subset of 5,200 active users.
For all real-world datasets, we use the pre-processing and splits from Zeng et al. (2024).
To compare against the other models, FIM-PP uses the sequences which the other models used for training as context and used the same inference sequences for evaluation.
Appendix F Evaluation Metrics
Following Zeng et al. (2024), we adopt a comprehensive set of metrics to evaluate both the temporal and categorical aspects of the predicted sequences. Let $\mathcal{S}_{\text{future}}=\{(t_{i},\kappa_{i})\}_{i=1}^{N}$ be a ground truth sequence of $N$ future events, and let $\hat{\mathcal{S}}_{\text{future}}$ be the corresponding predicted sequence. The metrics are defined based on the sequence of inter-arrival times $\mathbf{\Delta t}=[\Delta t_{1},\dots,\Delta t_{N}]$ (where $\Delta t_{i}=t_{i}-t_{i-1}$) and the sequence of marks.
Optimal Transport Distance (OTD)
We use the Optimal Transport Distance (OTD) to provide a holistic measure of similarity between the predicted and ground truth event sequences (Mei et al., 2019). OTD calculates the minimum cost required to transform the predicted sequence $\hat{\mathcal{S}}_{\text{future}}$ into the ground truth sequence $\mathcal{S}_{\text{future}}$ through a series of operations (insertions, deletions, and substitutions), each associated with a cost. This metric effectively captures discrepancies in timing, marks, and the total number of events.
RMSE on Event Counts (RMSEee)
This metric evaluates how well the model captures the distribution of event types in the predicted sequence. For each event type $\kappa\in\mathcal{K}$, we count its occurrences in the ground truth sequence ($C_{\kappa}$) and the predicted sequence ($\hat{C}_{\kappa}$). The RMSEee is the root mean squared error over the vector of these counts, averaged across all $m$ test sequences:
$\text{RMSE\textsubscript{$e$}}=\sqrt{\frac{1}{m}\sum_{j=1}^{m}\sum_{\kappa\in\mathcal{K}}(C_{j,\kappa}-\hat{C}_{j,\kappa})^{2}}$ (22)
Event Type Accuracy (Acc)
This metric directly measures the model’s ability to predict the correct event type at each position in the sequence. It is calculated as the fraction of events for which the predicted mark $\hat{\kappa}_{i}$ matches the ground truth mark $\kappa_{i}$, averaged over all test sequences. This provides a strict, position-wise evaluation of the categorical predictions.
$\text{Acc}=\frac{1}{m}\sum_{j=1}^{m}\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(\kappa_{j,i}=\hat{\kappa}_{j,i})$ (23)
where $\mathbb{I}(\cdot)$ is the indicator function. Unlike RMSEee, which assesses the overall distribution of event types, accuracy penalizes mispredictions at specific positions, making it a more challenging metric for sequential order. A higher accuracy indicates better performance.
Time-series Forecasting Metrics
To specifically assess the accuracy of the predicted inter-arrival times $\mathbf{\Delta t}$, we report two standard time-series forecasting metrics.
[…]
    $\text{RMSE\textsubscript{$\Delta t$}}=\sqrt{\frac{1}{m}\sum_{j=1}^{m}\frac{1}{N}\sum_{i=1}^{N}(\Delta t_{j,i}-\hat{\Delta}t_{j,i})^{2}}$ (24)
[…]
    $\text{sMAPE\textsubscript{$\Delta t$}}=\frac{100}{m}\sum_{j=1}^{m}\frac{1}{N}\sum_{i=1}^{N}\frac{2\lvert\Delta t_{j,i}-\hat{\Delta}t_{j,i}\rvert}{\lvert\Delta t_{j,i}\rvert+\lvert\hat{\Delta}t_{j,i}\rvert}$ (25)
Appendix G Challenges in Next Event Prediction
Our evaluations in table 2 reveal that FIM-PP in zero-shot mode already performs well for next event time prediction. It however gets a noticeably worse error in the next event type prediction. Up on investigating this, we found that many of the real-world datasets have specific patterns (such as oscillations between two marks) that FIM-PP (zs) struggles to capture (see fig 13). After fine-tuning, it is however able to spot these patterns well (see fig 14). This might also explain why FIM-PP performs better on long-horizon tasks: The specific order of the events does not matter here.
Figure 13: FIM-PP in zero-shot mode struggles to predict the next event type right if the dataset has alternating patterns such as here for the Taxi dataset.
Figure 14: After fine-tuning, FIM-PP is able to spot the alternating pattern between mark 3 and mark 8 in the Taxi dataset.
[…]
We hypothesize that the underlying reason for these two shortcomings is that our synthetic dataset distributions does not capture these patterns well. We are planning to investigate this further and to update our synthetic distribution to include such patterns and provide an updated version of FIM-PP.